### v1

# Projected and Solvable Topological Heavy Fermion Model of Twisted Bilayer Graphene

Abstract
We investigate the topological heavy-fermion (THF) model of magic-angle twisted bilayer graphene (MATBG) in the projected limit, where only the flat bands are present in the low-energy spectrum. Such limit has been previously analyzed in momentum-space Bistritzer-MacDonald-type continuum models, but not in a real-space formalism. In this regime, the Hubbard interaction ($U_{1}$) of the $f$-electrons is larger than the bandwidth ($2M$) of the flat bands but smaller than the gap ($\gamma$) between the flat and remote bands. In the THF model, concentrated charge (in real space) and concentrated Berry curvature (in momentum space) are respectively realized by exponentially localized $f$-orbitals and itinerant Dirac $c$-electrons. Local moments naturally arise from $f$-orbitals. Hybridizing the $f$-electrons with $c$-electrons produces power-law tails of the flat-band Wannier functions, raising the question of relevance of the local moment picture in the projected $U_{1}\ll\gamma$ limit. Nonetheless, we find that the local moments remain stable as long as $U_{1}\gg\Delta(\omega)$ for $|\omega|\lesssim U_{1}$, where $\Delta(\omega)\sim\gamma^{2}N(\omega)$ is the hybridization function seen by each $f$-site, and $N(\omega)$ is the density of states of the Dirac $c$-bands. Notably, the comparison between $U_{1}$ and $\gamma$ is irrelevant to the local moment formation if $N(\omega)$ is unknown. Within the framework of THF, we also derive the correlated self-energy of the flat bands using the Hubbard-I approximation, and estimate the coupling strength between the local moments. Finally, we comment that, in the regime of extremely concentrated Berry curvature, the single-particle gap between flat bands and remote bands vanishes and the interaction strength is always larger than the gap.
I Introduction and motivation
The emergence of flat bands in twisted bilayer graphene gives rise to a rich variety of strongly correlated phenomena Cao et al. (2018a); Kerelsky et al. (2019); Xie et al. (2019); Sharpe et al. (2019); Jiang et al. (2019); Choi et al. (2019); Polshyn et al. (2019); Yankowitz et al. (2019); Lu et al. (2019); Stepanov et al. (2020); Saito et al. (2020); Serlin et al. (2020); Chen et al. (2020); Wong et al. (2020a); Choi et al. (2020); Nuckolls et al. (2020); Choi et al. (2021a); Saito et al. (2021a); Liu et al. (2021a); Park et al. (2021); Wu et al. (2021a); Cao et al. (2021a); Das et al. (2021); Tschirhart et al. (2021); Pierce et al. (2021); Stepanov et al. (2021); Choi et al. (2021b); Xie et al. (2021a); Das et al. (2022); Nuckolls et al. (2023a); Yu et al. (2023a), including unconventional superconductivity Cao et al. (2018b); Yankowitz et al. (2019); Lu et al. (2019); Stepanov et al. (2020); Saito et al. (2020); de Vries et al. (2021); Oh et al. (2021a); Tian et al. (2023); Di Battista et al. (2022); Călugăru et al. (2022) and exotic quantum phases Tomarken et al. (2019); Cao et al. (2020); Zondiner et al. (2020a); Lisi et al. (2021); Benschop et al. (2021); Lian (2021); Rozen et al. (2021); Saito et al. (2021b); Lu et al. (2021); Hesp et al. (2021); Díez-Mérida et al. (2023); Hubmann et al. (2022); Ghawri et al. (2022); Jaoui et al. (2022); Paul et al. (2022); Grover et al. (2022); Zhou et al. (2023). Substantial theoretical efforts have been made to understand the rich physics of magic-angle twisted bilayer graphene (MATBG), focusing on constructing models Lopes dos Santos et al. (2007); Suárez Morell et al. (2010); Bistritzer and MacDonald (2011); Uchida et al. (2014); van Wijk et al. (2015); Dai et al. (2016); Jain et al. (2016); Nam and Koshino (2017); Efimkin and MacDonald (2018); Kang and Vafek (2018); Zou et al. (2018); Po et al. (2019); Liu et al. (2019a); Tarnopolsky et al. (2019); Mora et al. (2019); Li et al. (2019); Fang et al. (2019); Khalaf et al. (2019); Carr et al. (2019a, b); Rademaker et al. (2019); Kwan et al. (2020); Carr et al. (2020a); Tritsaris et al. (2020); Wilson et al. (2020); Park et al. (2020); Carr et al. (2020b); Fu et al. (2020); Huang et al. (2020); Calderón and Bascones (2020); Wu et al. (2021b); Ren et al. (2021); Hejazi et al. (2021); Călugăru et al. (2021); Bernevig et al. (2021a, b); Wang et al. (2021a); Ramires and Lado (2021); Shin et al. (2021); Lei et al. (2021); Cao et al. (2021b); Sheffer and Stern (2021); Wu et al. (2021c); Koshino et al. (2018); Ledwith et al. (2021); Guerci et al. (2022); Davydov et al. (2022); Classen et al. (2022); Lin et al. (2022); Samajdar et al. (2022); Kang and Vafek (2023); Vafek and Kang (2023); Shin et al. (2023), correlated states Ochi et al. (2018); Thomson et al. (2018); Xu et al. (2018); Koshino et al. (2018); Po et al. (2018); Venderbos and Fernandes (2018); Yuan and Fu (2018); Dodaro et al. (2018); Padhi et al. (2018); Kennes et al. (2018); Rademaker and Mellado (2018); Liu et al. (2019b); Huang et al. (2019); Wu et al. (2019a); Classen et al. (2019); Kang and Vafek (2019); Seo et al. (2019); Da Liao et al. (2019); Angeli et al. (2019); Xie and MacDonald (2020); Bultinck et al. (2020a); Chatterjee et al. (2020); Repellin et al. (2020); Cea and Guinea (2020); Zhang et al. (2020); Kang and Vafek (2020); Bultinck et al. (2020b); Chichinadze et al. (2020a); Soejima et al. (2020); Christos et al. (2020); Eugenio and Dag (2020); Wu and Das Sarma (2020); Vafek and Kang (2020); Xie et al. (2021b); Kang et al. (2021); Liu et al. (2021b); Da Liao et al. (2021); Liu and Dai (2021); Thomson and Alicea (2021); Kwan et al. (2021a); Lian et al. (2021); Zhang et al. (2021); Parker et al. (2021); Vafek and Kang (2021); Kwan et al. (2021b); Chen et al. (2021); Potasz et al. (2021); Xie et al. (2021c); Xie and MacDonald (2021); Ledwith et al. (2021); Cha et al. (2021); Kwan et al. (2021c); Hofmann et al. (2022); Wagner et al. (2022); Christos et al. (2022); Brillaux et al. (2022); Călugăru et al. (2022); Hong et al. (2022); Zhang et al. (2023); Blason and Fabrizio (2022); Xie et al. (2023); Kwan et al. (2023); Yu et al. (2023b); Wang et al. (2024a); Fernandes and Venderbos (2020); Datta et al. (2023a); Rai et al. (2024), and superconductivity Liu et al. (2024); Guo et al. (2018); Yuan and Fu (2018); Xu and Balents (2018); Dodaro et al. (2018); Po et al. (2018); Liu et al. (2018); Venderbos and Fernandes (2018); Isobe et al. (2018); Peltonen et al. (2018); Kennes et al. (2018); Wu et al. (2018); Guinea and Walet (2018); González and Stauber (2019); Huang et al. (2019); Roy and Juričić (2019); Wu et al. (2019b, a); You and Vishwanath (2019); Classen et al. (2019); Lian et al. (2019); Hu et al. (2019); Julku et al. (2020); Xie et al. (2020); Chichinadze et al. (2020b); Lopez-Bezanilla and Lado (2020); König et al. (2020); Christos et al. (2020); Wang et al. (2021b); Khalaf et al. (2021); Lewandowski et al. (2021); Fernandes and Fu (2021); Qin and MacDonald (2021); Phong et al. (2021); Choi and Choi (2021); Lake and Senthil (2021); Chou et al. (2021); Li et al. (2022); Fischer et al. (2022); Yu et al. (2022); Scammell et al. (2022); Christos et al. (2022); Chatterjee et al. (2022); Kwan et al. (2022); Wagner et al. (2023); González and Stauber (2023); Wagner et al. (2024); Wang et al. (2024b). The topological heavy-fermion (THF) model Song and Bernevig (2022); Călugăru et al. (2023); Shi and Dai (2022) has been proposed as a framework for comprehensively understanding MATBG. The THF model consists of $f$-orbitals, exponentially localized at the AA-stacking regions of the moiré pattern, and itinerant Dirac $c$-electrons. In this model, the strong correlations observed in MATBG are attributed to the local Coulomb repulsion of the quantum-dot-like $f$-orbitals, while the topology of the flat bands is carried by the Dirac $c$-electrons. The $cf$ hybridization generates a topological flat band that accurately reproduces both the dispersion and interaction form factors Călugăru et al. (2023) of the continuous Bistritzer-MacDonald (BM) model Bistritzer and MacDonald (2011). The local representation ($f$-orbitals) allows the use of advanced theoretical and numerical methods to study the correlation effect in MATBG Hu et al. (2023a); Zhou et al. (2024); Rai et al. (2023); Datta et al. (2023b); Song and Bernevig (2022); Shi and Dai (2022); Chou and Das Sarma (2023a); Lau and Coleman (2023); Wang et al. (2024c); Youn et al. (2024). Thus, various phases, including heavy Fermi liquid, disordered local moments arising from $f$-orbitals, RKKY-stabilized correlated insulators, unconventional superconductor and their competitions have been studied within this model using Hartree-Fock Song and Bernevig (2022); Shi and Dai (2022), large-$U$ perturbation theory Hu et al. (2023b), slave-particle Chou and Das Sarma (2023a); Lau and Coleman (2023), dynamical mean field theory (DMFT) Hu et al. (2023a); Zhou et al. (2024); Rai et al. (2023); Datta et al. (2023b); Youn et al. (2024), Gutzwiller method Hu et al. (2024), scaling theory Chou and Das Sarma (2023b); Zhou et al. (2024), and renormalized perturbation theory Wang et al. (2024c). Within a unified picture, these works provide explanations for experimental phenomena including quantum-dot-like cascade and zero-energy peak seen in STM spectrum, large entropy upon heating, sawtooth compressibility, V-shaped pairing gap, strong coupling feature of the superconductivity, and a link between cascade transitions and correlated Chern insulators.
THF model Ref. Ledwith et al. (2024) (non-local moments)
Wave function $e^{\mathrm{i\>\!}\frac{\pi}{4}l}\cdot\frac{1+\frac{(x+iy)/\lambda}{\mathrm{i\>\!}l(k_{x}+ik_{y})/(\gamma/v_{\star})}}{\sqrt{1+\frac{\gamma^{2}/v_{\star}^{2}}{\mathbf{k}^{2}}}}\cdot\frac{1}{\sqrt{2\pi}\lambda}\sum_{\mathbf{R}}e^{-\frac{|\mathbf{r}-\mathbf{R}|^{2}}{2\lambda^{2}}}$ $\chi_{l\beta}\cdot\frac{1+\frac{(x+iy)/\delta}{i(k_{x}+ik_{y})/s}}{\sqrt{1+\frac{2s^{2}}{\mathbf{k}^{2}}}}\cdot\frac{1}{\sqrt{2\pi}\delta}\sum_{\mathbf{R}}e^{-\frac{|\mathbf{r}-\mathbf{R}|^{2}}{4\delta^{2}}}$
Berry curvature $\Omega_{\mathbf{k}}=\frac{2\gamma^{2}/v_{\star}^{2}}{(\gamma^{2}/v_{\star}^{2}+|\mathbf{k}|^{2})^{2}}$ $\Omega_{\mathbf{k}}=\frac{4s^{2}}{(2s^{2}+|\mathbf{k}|^{2})^{2}}$
Self-energy at $\nu=0$ $\frac{U_{1}^{2}|\mathbf{k}|^{2}}{4i\omega(|\mathbf{k}|^{2}+\gamma^{2}/|v_{\star}|^{2})}$ (Hubbard-I approximation) $\frac{U_{1}^{2}|\mathbf{k}|^{2}}{4i\omega(|\mathbf{k}|^{2}+2s^{2})}$
Ferromagnetic coupling strength $\sim-U_{1}\frac{\gamma^{2}}{v_{\star}^{2}\Lambda_{c}^{2}}\log(\frac{|\gamma|}{|v_{\star}\Lambda_{c}|})$ $\sim-U_{\Gamma}s^{2}\log(s)$
Table 1: To establish the equivalence between our results and those in Ref. Ledwith et al. (2024), we let $s=|\gamma/(v_{\star}\sqrt{2})|$ and $\delta\equiv\lambda$. In addition, Ref. Ledwith et al. (2024) employs a different convention for the interaction strength. In the table above, we have converted their expressions to match the convention used in our work.
Previous work based on the THF model has primarily focused on the parameter region where the Hubbard repulsion ($U_{1}$) of $f$ electrons exceeds the gap ($\gamma$) between the remote bands and flat bands, i.e., $U_{1}>\gamma$. However, some worksLiu et al. (2024) have projected Hamiltonian (such as electron-phonon couplings), onto the flat bands using the wavefunction previously derived in Song and Bernevig (2022). In this study, we explore the THF model in the projected $U_{1}<\gamma$ limit, where the Hubbard repulsion of $f$ electrons is smaller than the gap between the flat bands and remote bands and larger than the bandwidth of flat bands. Previous studies of the projected limit have relied on the momentum-space Bistritzer-MacDonald model with projected Coulomb repulsion Kang and Vafek (2020); Kang et al. (2021); Vafek and Kang (2020); Kang and Vafek (2019); Xie et al. (2022); Kumar et al. (2021); Kwan et al. (2021c); Bernevig et al. (2021b); Lian et al. (2021); Bernevig et al. (2021c); Bultinck et al. (2020b); Potasz et al. (2021); Wu et al. (2018), where features of the “heavy particles”, such as the flat excitation spectrum near the Brillouin zone boundary, are clearly observed Kang et al. (2021); Bernevig et al. (2021b); Lian et al. (2021); Bernevig et al. (2021c); Bultinck et al. (2020b). These features seen in the momentum-space model are naturally explained by the THF model. However, a local description of only the flat bands is impossible due to their topological nature. The closest approximation is the nonlocal Wannier function of the topological in Ref. Ledwith et al. (2024), which we show shares many similarities with the THF framework, including the wavefunction initially obtained in Ref. Song and Bernevig (2022) (Eq. (S123)). In addition, we do not parametrically let the Berry curvature become delta function concentrated, as in that case the gap $\gamma$ would necessarily also vanish, which would invalidate the projection to the flat bands for any finite $U_{1}$.
In this work, we explore the correlation effect in the projected limit of the THF model. In Section II, we first review the THF model. In Section III, we explore both $k$-space and real-space non-local Wannier functions for topological flat bands within the THF model framework, as previously discussed in Ref. Song and Bernevig (2022). In Section IV, we explore the correlation effect in the projected limit of the THF. In Section V and in Table 1, we provide the comparison between the results obtained from THF model and the results in Ref. Ledwith et al. (2024). In Section VI, we discuss the realistic parameters of the THF model. Finally, we summarize our results in Section VII.
II The topological heavy fermion model
The single-particle Hamiltonian of THF model is
$\hat{H}_{0}=$ $\sum_{\eta s}\sum_{aa^{\prime}}\sum_{\mathbf{k},\mathbf{G}}(H^{(c,\eta)}_{a,a^{\prime}}(\mathbf{k}+\mathbf{G})-\mu\delta_{aa^{\prime}})c_{\mathbf{k}+\mathbf{G},a,\eta,s}^{\dagger}c_{\mathbf{k}+\mathbf{G},a^{\prime}\eta s}-\mu\sum_{\eta s}\sum_{\mathbf{R}}f_{\mathbf{R}\alpha\eta s}^{\dagger}f_{\mathbf{R}\alpha\eta s}$
$+\frac{1}{\sqrt{N_{M}}}\sum_{\eta s\alpha s}\sum_{\mathbf{R}}\sum_{\mathbf{k}\mathbf{G}}\left(e^{-\frac{\lambda}{2}|\mathbf{k}+\mathbf{G}|^{2}-i\mathbf{k}\cdot\mathbf{R}}H_{a\alpha}^{(cf,\eta)}(\mathbf{k})c_{\mathbf{k}+\mathbf{G},a,\eta,s}^{\dagger}f_{\mathbf{R}\alpha\eta s}+h.c.\right)\ .$ (1)
Here $\mathbf{R}$ sums over all the triangular moiré lattice formed by the AA-stacking regions of MATBG, $\mathbf{k}$ takes value in the moiré Brillouin zone, $\mathbf{G}$ sums over the moiré reciprocal lattice, and $N_{M}$ is the number of moiré unit cell. $\eta=\pm$ is the index for graphene valleys, $s=\uparrow,\downarrow$ is the spin index, $\alpha=1,2$ is the orbital index for $f$-electrons, and $a=1,2,3,4$ is the orbital index for $c$-electrons. The $f$-orbital can be expressed in terms of the continuous basis of the BM model
$f_{\mathbf{R},\alpha,\eta,s}^{\dagger}=\sum_{l\beta}\int d^{2}\mathbf{r}\ w_{l\beta,\alpha}^{(\eta)}(\mathbf{r}-\mathbf{R})c_{l,\beta,\eta,s}^{\dagger}(\mathbf{r})\ ,$ (2)
where $c_{l,\beta,\eta,s}^{\dagger}(\mathbf{r})$ creates a graphene electron at the (coarse-grained and continuous) position $\mathbf{r}$ in the valley $\eta$, layer $l$, sub-lattice $\beta$, spin $s$. Here $\beta=1,2$ corresponds to the A and B sublattices of graphene, respectively; $l=+,-$ corresponds to the top and bottom layer, respectively. $w_{l\beta,\alpha}^{(\eta)}(\mathbf{r}-\mathbf{R})$ is an exponentially localized Wannier function with a spread $\lambda$ about $1/2$ of the moiré lattice constant. The $c$-electrons can also be expressed in terms of the continuous basis as
$c_{\mathbf{k}+\mathbf{G},a,\eta,s}^{\dagger}=\frac{1}{\sqrt{N_{M}\Omega_{M}}}\sum_{l\beta}\int d^{2}\mathbf{r}\ \widetilde{u}^{(\eta)}_{l\beta,a}(\mathbf{r};\mathbf{k})c_{l\beta\eta s}^{\dagger}(\mathbf{r})\ ,$ (3)
where $\Omega_{M}$ is the area of the moiré unit cell. Ref. Song and Bernevig (2022) numerically constructed $\widetilde{u}^{(\eta)}_{l\beta,a}(\mathbf{r};\mathbf{k})$ around $\mathbf{k}=0$. Ref. Călugăru et al. (2023) (in its Appendix F, Eqs. F1-F4)) further provided a Gaussian form for it at momentum away from 0.
$\mu$ and $\lambda$ in Section II are the chemical potential and the spread of $f$-orbitals, respectively. The hopping matrices $H^{(c,\eta)}(\mathbf{k})$ and $H^{(cf,\eta)}(\mathbf{k})$ are
$H^{(c,\eta)}(\mathbf{k})=\begin{pmatrix}0_{2\times}&v_{\star}(\eta k_{x}\sigma_{0}+ik_{y}\sigma_{z})\\ v_{\star}(\eta k_{x}\sigma_{0}-ik_{y}\sigma_{z})&M\sigma_{x}\end{pmatrix},\qquad H^{(cf,\eta)}(\mathbf{k})=\begin{pmatrix}\gamma\sigma_{0}+v_{\star}^{\prime}(\eta k_{x}\sigma_{x}+k_{y}\sigma_{y})\\ 0_{2\times 2}\end{pmatrix}\ .$ (4)
The parameters $M$, $v_{\star}$, $\gamma$, $v_{\star}^{\prime}$, $\lambda$ have been extracted numerically and analytically from the BM model in a first-principle spirit Song and Bernevig (2022); Călugăru et al. (2023). Thus, they are functions of the twist angle $\theta$ and the ratio $w_{0}/w_{1}$ between inter-layer hoppings in the AA-stacking regions and AB-stacking regions. The two flat bands can hold eight electrons per moiré unit cell. We use $\nu$ to denote the number of electrons ($\nu>0$) or holes ($\nu<0$) per moiré unit cell counted from the charge neutrality point (CNP). $\nu=\pm 4$ corresponds to the band gaps between the flat bands and remote bands.
We do not give the explicit interaction Hamiltonian $H_{I}$ here. Interested readers may refer to Ref. Song and Bernevig (2022) for full details. The most important term in $H_{I}$ is an onsite Hubbard interaction $\hat{H}_{U_{1}}$
$\hat{H}_{U_{1}}=\frac{U_{1}}{2}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}:f_{\mathbf{R},\alpha,\eta,s}^{\dagger}f_{\mathbf{R},\alpha,\eta,s}::f_{\mathbf{R},\alpha^{\prime},\eta^{\prime},s^{\prime}}^{\dagger}f_{\mathbf{R},\alpha^{\prime},\eta^{\prime},s^{\prime}}:$ (5)
where $U_{1}$ is the interaction strength and $:f_{\mathbf{R},\alpha,\eta,s}^{\dagger}f_{\mathbf{R},\alpha,\eta,s}:=f_{\mathbf{R},\alpha,\eta,s}^{\dagger}f_{\mathbf{R},\alpha,\eta,s}-\frac{1}{2}$. Other terms include the density-density interactions between $f$- and $c$-electrons $\hat{H}_{W}$, density-density interactions between nearest neighbors of $f$-electrons $H_{U_{2}}$, a Hund’s exchange between $f$- and $c$-electrons $\hat{H}_{J}$, and Coulomb repulsions between $c$ electrons $\hat{H}_{V}$. All the interaction parameters are also determinate functions of the $\theta$ and $w_{0}/w_{1}$ Călugăru et al. (2023). The single-particle and interaction parameters at the magic angle are given in supplementary tables S4 and S6 of Ref. Song and Bernevig (2022). The full interacting theory, with terms beyond $U_{1}$ treated at the mean-field level, is detailed in an extensive DMFT-Iterative Perturbation Theory–Hubbard-I manuscript Călugăru and others .
III Wavefunction of the topological flat band
Via the THF model Song and Bernevig (2022), we could analytically derive the wavefunction for the topological flat bands. We consider the chiral-flat limit of the THF model by setting $M=0,v_{\star}^{\prime}=0$ in Eq. 4. Then, one can obtain exact flat Chern band solutions straightforwardly. The creation operator for Bloch state with Chern number $\zeta$ in valley $\eta$ is (Eq. (S123) of Ref. Song and Bernevig (2022))
$d_{\mathbf{k},\zeta,\eta,s}^{\dagger}=\frac{1}{\sqrt{\mathcal{N}_{\mathbf{k}}}}f_{\mathbf{k},1,\eta,s}^{\dagger}+\frac{1}{\sqrt{\mathcal{N}_{\mathbf{k}}}}\sum_{\mathbf{G}}\frac{\frac{\gamma}{v_{\star}}e^{-\frac{1}{2}\lambda^{2}|\mathbf{k}+\mathbf{G}|^{2}}}{\eta(k_{x}+G_{x})+i(-1)^{\zeta+1}(k_{y}+G_{y})}c_{\mathbf{k}+\mathbf{G},3,\eta,s}^{\dagger}\ ,$ (6)
where $\mathcal{N}_{\mathbf{k}}=1+\sum_{\mathbf{G}}\frac{(\gamma/v_{\star})^{2}}{|\mathbf{k}+\mathbf{G}|^{2}}e^{-\lambda^{2}|\mathbf{k}+\mathbf{G}|^{2}}$ is a normalization factor. To derive the analytical formula of topological flat-band wavefunction, we omit all $|\mathbf{G}|>0$ components. Using Eqs. 2 and 3, we obtain the following Bloch function with Chern number $+1$ in valley $+$
$\psi_{l\beta;\mathbf{k}}(\mathbf{r})=\sqrt{N_{M}}\langle 0|c_{l\beta+s}(\mathbf{r})d_{\mathbf{k},+,+,s}^{\dagger}|0\rangle=\frac{1}{\sqrt{\mathcal{N}_{\mathbf{k}}}}\sum_{\mathbf{R}}w^{(+)}_{l\beta,1}(\mathbf{r}-\mathbf{R})e^{\mathrm{i\>\!}\mathbf{k}\cdot\mathbf{R}}+\frac{1}{\sqrt{\mathcal{N}_{\mathbf{k}}}}\frac{\frac{\gamma}{v_{\star}}e^{-\frac{1}{2}\lambda^{2}\mathbf{k}^{2}}}{k_{x}+ik_{y}}\widetilde{u}_{l\beta,3}^{(+)}(\mathbf{r};\mathbf{k})\ .$ (7)
In the chiral limit, $w^{(+)}_{l\beta,1}(\mathbf{r})$ takes the form $w^{(+)}_{l\beta,1}(\mathbf{r})=\delta_{\beta 1}\frac{1}{\sqrt{2\pi}\lambda}e^{i\frac{\pi}{4}l-\mathbf{r}^{2}/(2\lambda^{2})}$ (see Eq. (S53) of Ref. Song and Bernevig (2022)). Appendix F of Ref. Călugăru et al. (2023) also gave the analytical expression for $\widetilde{u}$
$\widetilde{u}_{l\beta,3}^{(+)}(\mathbf{r};\mathbf{k})\approx\delta_{\beta 1}\frac{1}{\Omega_{M}}\sum_{\mathbf{G}}e^{-\mathrm{i\>\!}\frac{\pi}{4}l-\frac{1}{2}\mathbf{q}^{2}\lambda^{\prime 2}+\mathrm{i\>\!}(\mathbf{k}+\mathbf{G})\cdot\mathbf{r}}(\mathrm{i\>\!}G_{x}-G_{y})\sqrt{2\pi\lambda^{\prime 4}}\ ,\qquad(\text{for}\;|\mathbf{k}|a_{M}\ll 1)$ (8)
where $a_{M}$ is the moiré lattice constant. Here $\lambda^{\prime}$ characterizes the spread of the $c$-electrons. It is usually larger than $\lambda$ as $c$-electron is less localized. Inserting the identity $\frac{1}{\Omega_{M}}\sum_{\mathbf{G}}\delta(\mathbf{q}-\mathbf{G})=\frac{1}{(2\pi)^{2}}\sum_{\mathbf{R}}e^{i\mathbf{q}\cdot\mathbf{R}}$, we have
$\widetilde{u}_{l\beta,3}^{(+)}(\mathbf{r};\mathbf{k})=$ $\delta_{\beta 1}\sum_{\mathbf{R}}\int\frac{d^{2}\mathbf{q}}{(2\pi)^{2}}e^{-\mathrm{i\>\!}\frac{\pi}{4}l-\frac{1}{2}\mathbf{q}^{2}\lambda^{\prime 2}+\mathrm{i\>\!}(\mathbf{k}+\mathbf{q})\cdot\mathbf{r}-\mathbf{q}\cdot\mathbf{R}}(\mathrm{i\>\!}q_{x}-q_{y})\sqrt{2\pi\lambda^{\prime 4}}$
$=$ $\delta_{\beta 1}\frac{1}{\sqrt{2\pi}\lambda^{\prime 2}}\sum_{\mathbf{R}}e^{-\mathrm{i\>\!}\frac{\pi}{4}l+\mathrm{i\>\!}\mathbf{k}\cdot(\mathbf{r}-\mathbf{R})-\frac{(\mathbf{r}-\mathbf{R})^{2}}{2\lambda^{\prime 2}}}(x-R_{x}+i(y-R_{y}))\ ,\qquad(|\mathbf{k}|a_{M}\ll 1)\ .$ (9)
To simplify the expression, we take take $\lambda^{\prime}=\lambda$ which leads to a $\mathbf{k}$-independent Gaussian factor $e^{-\frac{|\mathbf{r}-\mathbf{R}|^{2}}{2\delta^{2}}}$. As $\lambda$ is smaller than $a_{M}$, there are $\mathbf{k}^{2}\lambda^{2}\ll 1$ for $\mathbf{k}$ in first Brillouin zone and $\lambda^{2}\ll\mathbf{R}^{2}$ for nonzero $\mathbf{R}$. Then we obtain the wave function in first Brillouin and first unit cell as
$\psi_{l\beta;\mathbf{k}}(\mathbf{r})=\delta_{\beta 1}e^{\mathrm{i\>\!}\frac{\pi}{4}l}\cdot\frac{1+\frac{(x+iy)/\lambda}{\mathrm{i\>\!}l(k_{x}+ik_{y})/(\gamma/v_{\star})}}{\sqrt{1+\frac{\gamma^{2}/v_{\star}^{2}}{\mathbf{k}^{2}}}}\cdot\frac{1}{\sqrt{2\pi}\lambda}\sum_{\mathbf{R}}e^{-\frac{|\mathbf{r}-\mathbf{R}|^{2}}{2\lambda^{2}}}\ .$ (10)
which leads to a concentrated charge distribution.
In addition, we calculate the Berry curvature of the flat band. Starting from the ansatz given in Eq. 6, we make two additional approximations: (1) we omit all $|\mathbf{G}|>0$ component; (2) we drop the exponential factor by setting $\lambda=0$. This results in the following simplified ansatz:
$d_{\mathbf{k},1,+,s}^{\dagger}\approx\frac{1}{\sqrt{1+\frac{\gamma^{2}}{|v_{\star}\mathbf{k}|^{2}}}}f_{\mathbf{k},1,+,s}^{\dagger}+\frac{1}{\sqrt{1+\frac{\gamma^{2}}{|v_{\star}\mathbf{k}|^{2}}}}\frac{\frac{\gamma}{v_{\star}}}{k_{x}+ik_{y}}c_{\mathbf{k}+\mathbf{G},3,+,s}^{\dagger}\ ,$ (11)
Using Eq. 11, we could obtain the following expression for the concentrated Berry curvature of the Chern $+1$ band in valley $+$
$\Omega_{\mathbf{k}}=\frac{2\gamma^{2}v_{\star}^{2}}{(\gamma^{2}+v_{\star}^{2}|\mathbf{k}|^{2})^{2}}\ .$ (12)
From Eqs. 10 and 12, we observe that our wavefunction of topological flat bands has two main features: concentrated charge in real space and concentrated Berry curvature in momentum space.
A similar wavefunction of the topological flat band has also been derived in Ref. Ledwith et al. (2024). The wavefunction derived in Ref. Ledwith et al. (2024) for Chern number $\zeta=1$ valley $\eta=+$ takes the form of
$\psi^{\rm NW}_{l\beta;\mathbf{k}}(\mathbf{r})=\chi_{l\beta}\cdot\frac{1+\frac{(x+iy)/\delta}{i(k_{x}+ik_{y})/s}}{\sqrt{1+\frac{2s^{2}}{\mathbf{k}^{2}}}}\cdot\frac{1}{\sqrt{2\pi}\delta}\sum_{\mathbf{R}}e^{-\frac{|\mathbf{r}-\mathbf{R}|^{2}}{4\delta^{2}}}\ ,$ (13)
where $\chi_{l\beta}$ is a spinor in valley and sub-lattice space, $\mathbf{r}\!=\!(x,y)$ is restricted in the first moiré unit cell, and $\mathbf{k}$ is restricted in the first moiré Brillouin zone. Here $\delta$ controls the spread of the concentrated charge in AA stacking regions, and $s$ controls the spread of Berry curvature in momentum space. When $\mathbf{r}$ ($\mathbf{k}$) exceeds the first unit cell (Brillouin zone), periodic conditions $\psi_{l\beta;\mathbf{k}+\mathbf{G}}^{\rm NW}(\mathbf{r})=\psi_{l\beta;\mathbf{k}}^{\rm NW}(\mathbf{r})$, $\psi_{l\beta;\mathbf{k}+\mathbf{G}}^{\rm NW}(\mathbf{r}+\mathbf{R})=e^{\mathrm{i\>\!}\mathbf{k}\cdot\mathbf{R}}\psi^{\rm NW}_{l\beta;\mathbf{k}}(\mathbf{r})$ are enforced. A subtlety of Eq. 13 is that $|\psi^{\rm NW}_{l\beta;\mathbf{k}}(\mathbf{r})|^{2}$ exhibits artificial discontinuities in both momentum and real spaces. In Ref. Ledwith et al. (2024) $\delta$ and $s$ are treated as fitting parameters to minimize the error in the intra-Chern-band form factors. The inter-Chern-band form factors are omitted. This automatically leads to an artificial chiral symmetry as well as a U(4)$\times$×U(4) symmetry in the projected interaction, as explained in Ref. Bernevig et al. (2021d), even though the parameters $\delta,s$ are fitted using realistic parameters without chiral symmetry ($w_{0}\neq 0$). We termed such projected interaction as in a quasi-chiral limit which has also been discussed in Ref. Song and Bernevig (2022).
Comparing Eq. 10 to Eq. 13, we identify the correspondence between THF parameters and NW ansatz parameters:
$s\equiv|\frac{\gamma}{\sqrt{2}v_{\star}}|,\qquad\delta\equiv\lambda\ .$ (14)
Our form in Eq. 10 obtained in the chiral-flat limit differs from the NW ansatz (Eq. 13) by a $\mathrm{i\>\!}l$ phase factor. Through the THF model, we could also go beyond the chiral-flat limit by introducing non-zero $M$ and $v_{\star}^{\prime}$ and obtain the flat-band wave function either numerically or perturbatively. Additionally, by including $c$-electrons for all the $\mathbf{G}$ shell in Eq. 6, our derived expressions Song and Bernevig (2022) for the wavefunction of flat bands do not exhibit any discontinuities. Eq. 10 (as also derived in Ref. Ledwith et al. (2024)) can be understood as a special case of it.
IV Correlation effects in the projected limit of the topological heavy-fermion model
In this section, we study the correlation effects of the heavy-fermion model in the projected limit
$|v_{\star}\Lambda_{c}|\gg|\gamma|\gg U_{1}$ (15)
and demonstrate the local moment formation in this limit of the THF model. Here, $\Lambda_{c}$ is the momentum cutoff of the $c$ electron, which is on the order of $\sim\frac{1}{2}|\mathbf{b}_{M,1}|$. Thus, $|v_{\star}\Lambda_{c}|$ can be interpreted as the bandwidth of $c$ electrons. In the projected limit, the gap between the flat band and remote band, $\gamma$, is larger than the Hubbard interaction $U_{1}$. Only the flat bands remain active and experience strong interaction effects. Moreover, since $\gamma/|v_{\star}\Lambda_{c}|\ll 1$, we only expect a non-zero Berry curvature in a small (but non-vanishing) region near $\Gamma$ point, where $|\mathbf{k}|<\frac{|v_{\star}|}{\gamma}$ (see Eq. 12). In other words, the band remains trivial over most of the moiré Brillouin zone. In addition, the orbital weight of $f$ electrons in the flat band is (from Eq. 11)
$\frac{1}{N_{M}}\sum_{\mathbf{k}}\frac{1}{1+\frac{\gamma^{2}}{|v_{\star}\mathbf{k}|^{2}}}\approx\frac{1}{\Omega_{M}}\int_{|\mathbf{k}|<\Lambda_{c}}\frac{1}{1+\frac{\gamma^{2}}{|v_{\star}\mathbf{k}|^{2}}}$ (16)
with $\Omega_{M}$ the size of the moiré Brillouin zone. In practice, we can set $\Omega_{M}\approx\pi\Lambda_{c}^{2}$. In other words, we approximate the original hexagonal Brillouin zone with a circular one. This leads to the following orbital weights for the $f$ electrons:
$\frac{1}{N_{M}}\sum_{\mathbf{k}}\frac{1}{1+\frac{\gamma^{2}}{|v_{\star}\mathbf{k}|^{2}}}\approx 1+\frac{\gamma^{2}}{|v_{\star}\Lambda_{c}|^{2}}\log\frac{1}{1+\frac{|v_{\star}\Lambda_{c}|^{2}}{\gamma^{2}}}$ (17)
For $|\gamma|/|v_{\star}\Lambda_{c}|\ll 1$, the orbital weight of $f$ electrons is $\sim 1-2\bigg{[}|\gamma|/|v_{\star}\Lambda_{c}|\bigg{]}^{2}\log(|\gamma/|v_{\star}\Lambda_{c}|)$, indicating that the flat bands are mostly formed by $f$ electrons in this limit. Since, in the projected limit, only the flat bands are the relevant low-energy degrees of freedom and these bands are primarily composed of $f$ electrons, we expect the correlation effects to be well described by the heavy-fermion $f$-electron basis and largely governed by the Hubbard interactions between $f$ electrons ($\hat{H}_{U_{1}}$).
In addition, the local moment formation can be understood by viewing each $f$-site as an Anderson impurity coupled to an effective bath describing its environment. Ref. Zhou et al. (2024) (Eq. (B25)) derived the hybridization function of the effective Anderson impurity in the flat-band limit ($M=0$)
$\Delta(\omega)=\frac{\sqrt{3}}{8}\cdot\frac{a_{M}^{2}\gamma^{2}}{v_{\star}^{2}}|\omega|\ ,$ (18)
where $a_{M}\sim\Lambda_{c}^{-1}$ is the moiré lattice constant. As revealed in many calculations, the bandwidth of the renormalized spectrum is driven by $U_{1}$ (but can be renormalized down, to for example $0.6U_{1}$ in DMFT results Rai et al. (2023)). Thus, the typical hybridization within this energy scale, which characterizes the inversion lifetime of $f$-electrons, is given by $\Delta_{0}(U_{1})$. A local moment can be stabilized as long as
$U_{1}\gg\Delta_{0}(U_{1})\qquad\Leftrightarrow\qquad\frac{a_{M}\gamma}{v_{\star}}\ll 1\quad\text{or}\quad\frac{\gamma}{v_{\star}\Lambda_{c}}\ll 1\ .$ (19)
The Kondo temperature $T_{\rm K}\propto\exp(-\frac{\pi U_{1}}{32\Delta})$ is exponentially suppressed in this limit. Since $\frac{\gamma}{v_{\star}\Lambda_{c}}\ll 1$ implies $\gamma<100meV$, local moment formation happens in TBG.
In this section, we discuss the correlation effects in the projected limit (Eq. 15). We derive the self-energy of the flat bands, and also the effective coupling between flavor moments formed by the flat bands. To simplify the analysis, we will take the chiral flat limit with $v_{\star}^{\prime}=0,M=0$. Since the flat bands are primarily composed of $f$ electrons, we expect the correlation effects to be driven by the Hubbard interaction of the $f$ electrons. Therefore, we only treat $H_{W},H_{V}$ at the Hartree-Fock level, where they contribute only single-particle terms to the Hamiltonian. We will also drop the $H_{J}$ which is generally much weaker than the other interactions.
IV.1 Self-energy and Hubbard-I approximation
To obtain the self-energy, we first integrate out the $c$ electrons, leading to the following effective action of $f$ electrons
$S_{f}=$ $\frac{1}{\beta}\sum_{\mathbf{R},\alpha\eta s,\omega}f_{\mathbf{R},\alpha\eta s}^{\dagger}(i\omega)\big{(}i\omega-\Sigma_{0}(\mathbf{R}-\mathbf{R}^{\prime},\mathrm{i\>\!}\omega)+\epsilon_{f}\big{)}f_{\mathbf{R},\alpha\eta s}(i\omega)$
[…]
Here, the $\Sigma_{0}$ term denotes the contribution from integrating out $c$ electrons, $\epsilon_{f}$ is the chemical potential term of $f$ electrons which includes the Hartree contributions from $H_{U_{2}},H_{W}$ Hu et al. (2023b), $\tau$ is imaginary time, $\omega$ is Matsubara frequency, and $\beta$ is the inverse temperature. Explicitly, $\Sigma_{0}$ is given by
$\Sigma_{0}(\mathbf{R}-\mathbf{R}^{\prime},i\omega)=$ $\frac{1}{\Omega_{M}}\int_{|\mathbf{k}|<\Lambda_{c}}\frac{d^{2}\mathbf{k}}{(2\pi)^{2}}\bigg{[}\frac{\gamma^{2}/2}{\mathrm{i\>\!}\omega-|v_{\star}\mathbf{k}|-\epsilon_{c}}+\frac{\gamma^{2}/2}{\mathrm{i\>\!}\omega+|v_{\star}\mathbf{k}|-\epsilon_{c}}\bigg{]}e^{i\mathbf{k}\cdot(\mathbf{R}-\mathbf{R}^{\prime})}$
[…]
where $\epsilon_{c}$ describes the Hartree contributions to the $c$ electrons from the $\hat{H}_{W},\hat{H}_{V}$ Călugăru and others ; Rai et al. (2023); Hu et al. (2023b). From Eq. 21, we observe the contribution from $c$ electrons approaches zero as $\frac{\gamma^{2}}{v_{\star}^{2}\Lambda_{c}^{2}}\rightarrow 0$. However, it is worth mentioning that the exact limit $\frac{\gamma^{2}}{v_{\star}^{2}\Lambda_{c}^{2}}=0$ also indicates the gap vanishes ($\gamma\rightarrow 0$), given that $v_{\star},\Lambda_{c}$ are finite.
We first derive the self-energy of $f$ electrons in the limit of $|\gamma|/|v_{\star}\Lambda_{c}|\rightarrow 0$. In this limit, the effective action of $f$ electrons is
$S_{at}=\frac{1}{\beta}\sum_{\mathbf{R},\alpha\eta s,i\omega}f_{\mathbf{R},\alpha\eta s}^{\dagger}(i\omega)(i\omega-\epsilon_{f})f_{\mathbf{R},\alpha\eta s}(i\omega)+\frac{U_{1}}{2}\int_{\tau}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}:f_{\mathbf{R},\alpha\eta s}^{\dagger}(\tau):f_{\mathbf{R},\alpha\eta s}(\tau)::f_{\mathbf{R},\alpha^{\prime}\eta^{\prime}s^{\prime}}^{\dagger}(\tau):f_{\mathbf{R},\alpha^{\prime}\eta^{\prime}s^{\prime}}(\tau):$ (22)
This describes an atomic problem where $f$ electrons of different sites are decoupled. We use $\nu_{f}$ to characterize the filling of the $f$ electrons. The atomic action $S_{at}$ can be solved exactly which yields the following self-energy in the low-energy limit for the $f$ electrons at the integer filling ($1/\beta\ll U_{1}$ălugăru and others
$\Sigma_{f}(i\omega)=\frac{1}{64}\frac{U_{1}^{2}(4-\nu_{f})(\nu_{f}+4)}{i\omega-U_{1}\nu_{f}/8}$ (23)
Moreover, in the projected limit, the difference between the filling of $f$ electrons, $\nu_{f}$, and the total filling of the systems, $\nu$, is of the order $\gamma^{2}/|v_{\star}\Lambda_{c}|^{2}\log(|\gamma|/|v_{\star}\Lambda_{c}|)$, as estimated from the orbital weights of $f$ electrons of the flat bands (Eq. 17). Thus, as $|\gamma|/|v_{\star}\Lambda_{c}|\rightarrow 0$, the self-energy of $f$ electrons can be approximately written as
$\Sigma_{f}(i\omega)\approx\frac{1}{64}\frac{U_{1}^{2}(4-\nu)(\nu+4)}{i\omega-U_{1}\nu/8}$ (24)
Using Eq. 24, we can now investigate the single-particle excitation by evaluating the interacting Green’s functions. We define the full Green’s function as
$G^{\alpha\eta s}(\tau,\mathbf{k})=\begin{bmatrix}-\langle T_{\tau}f_{\mathbf{k},\alpha\eta s}(\tau)f_{\mathbf{k},\alpha\eta s}^{\dagger}\rangle&-\langle T_{\tau}f_{\mathbf{k},\alpha\eta s}(\tau)c_{\mathbf{k},\alpha\eta s}^{\dagger}\rangle&-\langle T_{\tau}f_{\mathbf{k},\alpha\eta s}(\tau)c_{\mathbf{k},\alpha+2\eta s}^{\dagger}\rangle\\ -\langle T_{\tau}c_{\mathbf{k},\alpha\eta s}(\tau)f_{\mathbf{k},\alpha\eta s}^{\dagger}\rangle&-\langle T_{\tau}c_{\mathbf{k},\alpha\eta s}(\tau)c_{\mathbf{k},\alpha\eta s}^{\dagger}\rangle&-\langle T_{\tau}c_{\mathbf{k},\alpha\eta s}(\tau)c_{\mathbf{k},\alpha+2\eta s}^{\dagger}\rangle\\ -\langle T_{\tau}c_{\mathbf{k},\alpha+2\eta s}(\tau)f_{\mathbf{k},\alpha\eta s}^{\dagger}\rangle&-\langle T_{\tau}c_{\mathbf{k},\alpha+2\eta s}(\tau)c_{\mathbf{k},\alpha\eta s}^{\dagger}\rangle&-\langle T_{\tau}c_{\mathbf{k},\alpha+2\eta s}(\tau)c_{\mathbf{k},\alpha+2\eta s}^{\dagger}\rangle\end{bmatrix}$ (25)
Using the self-energy in Eq. 24, we obtain
$[G^{\alpha\eta s}(i\omega,\mathbf{k})]^{-1}=\begin{bmatrix}i\omega-\epsilon_{f}&-\gamma&0\\ -\gamma&i\omega-\epsilon_{c}&-v_{\star}(\eta k_{x}+i(-1)^{\alpha+1}k_{y})\\ 0&-v_{\star}(\eta k_{x}-i(-1)^{\alpha+1}k_{y})&i\omega-\epsilon_{c}\end{bmatrix}-\begin{bmatrix}\frac{1}{64}\frac{U_{1}^{2}(4-\nu)(\nu+4)}{i\omega-U_{1}\nu/8}&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix}$ (26)
The Green’s function in the above equation with atomic self-energy is equivalent to the Green’s function obtained from Hubbard-I approximation Hubbard (1963). We can determine the Hartree contributions $\epsilon_{f},\epsilon_{c}$ self-consistently at each filling, and then numerically obtain the spectral functions $\rho(\epsilon)=\frac{1}{\pi}\text{Tr}[G^{\alpha\eta s}(\omega-i0^{+})]$ (Fig. 1). We show the resulting spectral functions at $\nu=0,-1,-2,-3$ in Fig. 1, which is consistent with the result derived in Ref. Ledwith et al. (2024) via the momentum-space model.
We now note that, via the THF model, we can also evaluate the excitation spectrum analytically. Since we are in the limit of $\gamma\gg U_{1}$, we can project the Green’s function to the flat band. The wavefunctions of the flat bands are (see also Eq. 11)
$v^{\alpha\eta s}_{\mathbf{k}}=\frac{1}{\sqrt{1+\gamma^{2}/|v_{\star}\bf{k}|^{2}}}\begin{bmatrix}1&0&\frac{\gamma}{v_{\star}(\eta k_{x}+i(-1)^{\alpha+1}k_{y})}\end{bmatrix}^{T}$ (27)
It is worth mentioning that, away from CNP, non-zero $\epsilon_{f}$ and $\epsilon_{c}$ appear, which could modify the wave function of flat bands in Eq. 27. However, since $\epsilon_{f},\epsilon_{c}$ are both proportional to the interaction strength and are much smaller than $\gamma$ under our assumptions. In  Hubbard (1963) we give the expressions for the Hubbard-I for all range of parameters $\gamma,U_{1}$. We expect the flat-band wavefunction can still be approximately written as Eq. 27. To get a very simple analytical formula, we approximately set $\epsilon_{f}=\epsilon_{c}=0$ and project the Green’s function to the flat bands. This gives the following interacting Green’s function of flat bands
$[G_{\rm flat}^{\alpha\eta s}(i\omega,\mathbf{k})]^{-1}=v_{\mathbf{k}}^{\alpha\eta s,{\dagger}}\cdot[G^{\alpha\eta s}(\omega,\mathbf{k})]^{-1}\cdot v_{\mathbf{k}}^{\alpha\eta s}=i\omega-\frac{U_{1}^{2}|v_{\star}\mathbf{k}|^{2}(4-\nu)(4+\nu)}{64(i\omega-U_{1}\nu/8)(|v_{\star}\mathbf{k}|^{2}+\gamma^{2})}$ (28)
This also indicates the self-energy of the flat-bands can be written as
$\Sigma_{flat}(i\omega)=\frac{U_{1}^{2}|v_{\star}\mathbf{k}|^{2}(4-\nu)(4+\nu)}{64(i\omega-U_{1}\nu/8)(|v_{\star}\mathbf{k}|^{2}+\gamma^{2})}$ (29)
We could focus on $\nu=0$, the self-energy then becomes
$\Sigma_{flat}(i\omega)=\frac{U_{1}^{2}|v_{\star}\mathbf{k}|^{2}}{4i\omega(|v_{\star}\mathbf{k}|^{2}+\gamma^{2})}$ (30)
The single-particle excitation of the system can be obtained by finding the poles of Green’s function ($[G^{\alpha\eta s}_{flat}(i\omega\rightarrow E_{\mathbf{k}},\mathbf{k})]^{-1}=0$) which gives (at $\nu=0$)
$E_{\mathbf{k},1/2}=\pm\frac{1}{2}\sqrt{\frac{U_{1}^{2}|v_{\star}\mathbf{k}|^{2}}{|v_{\star}\mathbf{k}|^{2}+\gamma^{2}}}$ (31)
Away from $\Gamma$ point with $|v_{\star}\mathbf{k}|\gg\gamma$, the single-particle excitation shows Mott-like behavior and are gapped with $E_{\mathbf{k},1/2}\approx\pm\frac{1}{2}U_{1}$. Near the $\Gamma$ point with $|v_{\star}\mathbf{k}|\ll\gamma$, we obtain a Dirac dispersion with
$E_{\mathbf{k},1/2}\approx\pm\frac{U_{1}}{2|\gamma|}|v_{\star}\mathbf{k}|$ (32)
At $\nu\neq 0$, due to the absence of particle-hole symmetry, the Dirac nodes will be gapped (Fig. 1).
Finally, the effective action $S_{at}$ (Eq. 22) describes an atomic problem with Hubbard interactions. The entropy of this atomic problem can be solved exactly. At filling $\nu$, the number of $f$ electrons are $\nu_{f}+4\approx\nu+4$. These $f$ electrons can occupy $8(=2\times 2\times 2)$-degenerate $f$ orbitals. The ground states thus have a degeneracy of $\frac{8!}{(\nu+4)!(4-\nu)!}$. This gives the following entropy of the entire (unstrained) system
$S=k_{B}N_{M}\log(\frac{8!}{(\nu+4)!(4-\nu)!})$ (33)
The entropy of the THF with strain is solved in Zhang et al. .
Finally, we comment that we expect the Hubbard-I approximation to provide a qualitative good description of the single-particle dispersion, for both $U_{1}>|\gamma|$ and $U_{1}<|\gamma|$, as long as the system does not develop the Kondo effect or long-range ordering. In the low-temperature limit, long-range orders and the Kondo effect become important and can lead to a qualitative change of the self-energy behaviors as shown in the full DMFT solution Rai et al. (2023).
Figure 1: Spectrum obtained from Hubbard-I approximation at integer fillings $\nu=0,-1,-2,-3$ (from left to right). To realize the limit of $|v_{\star}\Lambda_{c}|\gg\gamma\gg U_{1}$, we have taken the parameter values given in Ref. Song and Bernevig (2022), and then rescaled all the interaction strength by a factor of $1/3$.
IV.1.1 Single-particle spectrum in the Hubbard-I approximation across all interaction regimes
In this section, we discuss the single-particle spectrum in the Hubbard-I approximation in detail. We first utilize the idea introduced in Section VII.1 and map the interacting Green’s function to an effective non-interacting Green’s function. We note that the self-energy in Eq. 24 can be written as
[…]
Therefore, we can introduce the following non-interacting systems and the auxiliary fermion operators $a$ (similarly as  Eq. 94)
$\hat{H}^{aux}=$ $\sum_{\mathbf{k},\alpha\eta s}[\psi_{\mathbf{k},\alpha\eta s}^{\dagger}]_{i}h_{ij}^{\alpha\eta s}(\mathbf{k})[\psi_{\mathbf{k},\alpha\eta s}]_{j}$
[…]
where
$\psi_{\mathbf{k},\alpha\eta s}=\begin{bmatrix}f_{\mathbf{k},\alpha\eta s}&c_{\mathbf{k},\alpha\eta s}&c_{\mathbf{k},(\alpha+2)\eta s}\end{bmatrix}$
$h^{\alpha\eta s}(\mathbf{k})=\begin{bmatrix}\epsilon_{f}&\gamma&0\\ \gamma&\epsilon_{c}&v_{\star}(\eta k_{x}+i(-1)^{\alpha+1}k_{y})\\ 0&v_{\star}(\eta k_{x}-i(-1)^{\alpha+1}k_{y})&\epsilon_{c}\end{bmatrix}$ (36)
Eq. 35 gives exactly the same Green’s function (for $f,c$ electrons) and the same single-particle spectrum as the Hubbard-I Green’s function  Eq. 26. Therefore, it is sufficient to analyze the single-particle spectrum of $\hat{H}^{aux}$ which can be described by the following single-particle Hamiltonian
$h^{aux,\alpha\eta s}(\mathbf{k})=\begin{bmatrix}\epsilon_{f}&\gamma&0&\frac{U_{1}\sqrt{(4-\nu_{f})(4+\nu_{f})}}{8}\\ \gamma&\epsilon_{c}&v_{\star}(\eta k_{x}+i(-1)^{\alpha+1}k_{y})&0\\ 0&v_{\star}(\eta k_{x}-i(-1)^{\alpha+1}k_{y})&\epsilon_{c}&0\\ \frac{U_{1}\sqrt{(4-\nu_{f})(4+\nu_{f})}}{8}&0&0&U_{1}\nu_{f}/8\end{bmatrix}$ (37)
This is true in all regimes of $U_{1}$. Our mapping to the effective system with auxiliary fermions reproduces the original Green’s function, which indicates the mapping preserves the symmetries. Since $f_{\mathbf{k},\alpha\eta s}$ and $a_{\mathbf{k},\alpha\eta s}$ are hybridized with a momentum independent hybridization term, which indicates $a_{\mathbf{k},\alpha\eta s}$ and $f_{\mathbf{k},\alpha\eta s}$ shares the same symmetry properties.
We first investigate the spectrum near $\Gamma$ point. The spectrum could be obtained by diagonalizing $H^{aux,\alpha\eta s}(\mathbf{k}=0)$. In the limit of $U_{1}\ll\gamma$, the $f_{\alpha\eta s},c_{\mathbf{k},\alpha\eta s}$ are high energy degrees of freedom and are gapped by large $|\gamma|$. The low-energy bands are described by $c_{\mathbf{k},(\alpha+2)\eta s},a_{\mathbf{k},\alpha\eta s}$. We can then treat the effect of $f_{\alpha\eta s},c_{\mathbf{k},\alpha\eta s}$ perturbatively by integrating them out. This then leads to the following effective Hamilton
$H^{eff}=\sum_{\mathbf{k},\alpha\eta s}\begin{bmatrix}c_{\mathbf{k},(\alpha+2)\eta s}^{\dagger}&a_{\mathbf{k},\alpha\eta s}^{\dagger}\end{bmatrix}h^{eff,\alpha\eta s}(\mathbf{k})\begin{bmatrix}c_{\mathbf{k},(\alpha+2)\eta s}\\ a_{\mathbf{k},\alpha\eta s}\end{bmatrix}$
$h^{eff,\alpha\eta s}(\mathbf{k})=\begin{bmatrix}\epsilon_{c}&\\ &\frac{U_{1}\nu_{f}}{8}\\ \end{bmatrix}+\frac{-U_{1}v_{\star}\sqrt{(4-\nu_{f})(4+\nu_{f})}}{8\gamma}\begin{bmatrix}0&\eta k_{x}-i(-1)^{\alpha+1}k_{y}\\ \eta k_{x}+i(-1)^{\alpha+1}k_{y}&0\end{bmatrix}$ (38)
The corresponding spectrum is
$\frac{8\epsilon_{c}+U_{1}\nu_{f}}{16}\pm\sqrt{\bigg{[}\frac{8\epsilon_{c}-U_{1}\nu_{f}}{8}\bigg{]}^{2}+\frac{U_{1}^{2}v_{\star}^{2}(4-\nu_{f})(4+\nu_{f})}{16^{2}\gamma^{2}}|\mathbf{k}|^{2}}$ (39)
We can observe that at charge neutrality, the system is gapless. However, away from charge neutrality, due to the effect of particle-hole symmetry breaking, the system is gapped.
We could further discuss the single-particle spectrum away from the flat band limit by including a finite $M$ term. Our effective Hamiltonian now becomes
$H^{eff,M}=\sum_{\mathbf{k},\alpha\eta s}\begin{bmatrix}c_{\mathbf{k},3\eta s}^{\dagger}&a_{\mathbf{k},1\eta s}^{\dagger}&c_{\mathbf{k},4\eta s}^{\dagger}&a_{\mathbf{k},2\eta s}^{\dagger}\end{bmatrix}h^{eff,\alpha\eta s}(\mathbf{k})\begin{bmatrix}c_{\mathbf{k},3\eta s}\\ a_{\mathbf{k},1\eta s}\\ c_{\mathbf{k},4\eta s}\\ a_{\mathbf{k},2\eta s}\end{bmatrix}$
$h^{eff,\alpha\eta s}(\mathbf{k})=\begin{bmatrix}\epsilon_{c}&0&M&0\\ 0&\frac{U_{1}\nu_{f}}{8}&0&0\\ M&0&\epsilon_{c}&0\\ 0&0&0&\frac{U_{1}\nu_{f}}{8}\end{bmatrix}+\frac{-U_{1}v_{\star}\sqrt{(4-\nu_{f})(4+\nu_{f})}}{8\gamma}\begin{bmatrix}0&\eta k_{x}-ik_{y}&0&0\\ \eta k_{x}+ik_{y}&0&0&0\\ 0&0&0&\eta k_{x}+ik_{y}\\ 0&0&\eta k_{x}-ik_{y}&0\end{bmatrix}$ (40)
The eigenvalues can be obtained exactly which are
$\frac{\epsilon_{c}+U_{1}\nu_{f}/8}{2}+\frac{M}{2}\pm\sqrt{\tilde{V}^{2}|\mathbf{k}|^{2}+\frac{1}{4}\bigg{(}\epsilon_{c}-U_{1}\nu_{f}/8+M\bigg{)}^{2}}$
$\frac{\epsilon_{c}+U_{1}\nu_{f}/8}{2}-\frac{M}{2}\pm\sqrt{\tilde{V}^{2}|\mathbf{k}|^{2}+\frac{1}{4}\bigg{(}\epsilon_{c}-U_{1}\nu_{f}/8-M\bigg{)}^{2}}$
where
$\tilde{V}=\frac{U_{1}\sqrt{(4-\nu_{f})(4+\nu_{4})}}{8\gamma}v_{\star}\,$ (42)
We now discuss the evolution of Dirac nodes. At charge neutrality with $\nu_{f}=0,\epsilon_{c}=0$, the spectrum becomes
±$\pm\bigg{(}\frac{M}{2}\pm\sqrt{\tilde{V}^{2}|\mathbf{k}|^{2}+M^{2}/4}\bigg{)}$ (43)
which leads to a quadratic Dirac node.
Away from charge neutrality, the spectrum at $\mathbf{k}=0$ is
$\epsilon_{c}+M,\quad\epsilon_{c}-M,\quad\frac{U_{1}\nu_{f}}{8},\quad\frac{U_{1}\nu_{f}}{8}$ (44)
The two-fold degenerate states give a quadratic Dirac node. However, the energy of the Dirac node is no longer at $E=0$ due to the particle-hole symmetry breaking at $\nu\neq 0$. Therefore, depending on the parameters of the model, we could have either a gap state or a gapless state. In practice, if
$\frac{U_{1}}{8}\nu_{f}>\epsilon_{c}+M>\epsilon_{c}-M$
or $\frac{U_{1}}{8}\nu_{f}<\epsilon_{c}+M>\epsilon_{c}-M\,,$ (45)
then the quadratic Dirac node is formed by the top two bands or the bottom two bands. The system can still develop a charge excitation gap at nonzero integer filling without breaking any valley or spin symmetries. This is also what we observed in our numerical calculations. Additionally, we find that no other nodes exist at other momentum points. On the other hand, if Section IV.1.1 is not satisfied, the system must be gapless since the middle two bands are degenerate at $\Gamma$ point.
IV.1.2 Comparison of the Hubbard-I spectrum in two limits
We now compare the Hubbard-I spectrum (at $M=0,v_{\star}^{\prime}=0$) in two limits: $U_{1}<\gamma$ and $U_{1}>\gamma$. Our focus is on the Fermi velocity of the gapless node at the $\Gamma$ point for $\nu=0$. At $\nu=0$, we have $\epsilon_{f}=0,\nu_{f}=0,epsilon_{c}=0$,. By diagonalizing the auxiliary Hamiltonian Eq. 37 within the Hubbard-I approximation, we obtain the following two eigenvalues that describe a Dirac node
±$\pm\sqrt{\frac{\gamma^{2}+|v_{\star}\mathbf{k}|^{2}+(\frac{U_{1}}{2})^{2}}{2}-\sqrt{\bigg{[}\frac{\gamma^{2}+|v_{\star}\mathbf{k}|^{2}+(\frac{U_{1}}{2})^{2}}{2}\bigg{]}^{2}-\frac{U_{1}^{2}|v_{\star}\mathbf{k}|^{2}}{4}}}$ (46)
A small momentum expansion leads to
±$\pm\sqrt{\frac{U_{1}^{2}}{4\gamma^{2}+U_{1}^{2}}}|v_{\star}\mathbf{k}|$ (47)
with velocity $v_{D}=\sqrt{\frac{U_{1}^{2}}{4\gamma^{2}+U_{1}^{2}}}v_{\star}$. We could then get the velocity in two limits which are
$U_{1}\gg|\gamma|$ $:\quad v_{D}\approx v_{\star}$
$U_{1}\ll|\gamma|$ $:\quad v_{D}\approx\frac{U_{1}}{2|\gamma|}v_{\star}$ (48)
In the limit of $U_{1}\gg|\gamma|$, the fermion velocity approaches that of graphene. As $U_{1}$ decreases, the fermion velocity becomes gradually renormalized. Taking the parameters derived from ab-initio and given in Ref. Song and Bernevig (2022), we observe $\sqrt{\frac{U_{1}^{2}}{4\gamma^{2}+U_{1}^{2}}}\approx 0.76$. Then the velocity of Dirac dispersion is $v_{D}\approx 3.27\text{eV}\cdot\text{\AA}$. However, we also comment that the effective Hubbard $U_{1}$ (which characterizes the separation of Hubbard bands) could be further renormalized due to the finite hybridization as shown in the DMFT simulation Rai et al. (2023). This could further renormalize the velocity of Dirac dispersions.
IV.2 Effective coupling between flavor moments
In the limit of $|\gamma|/|v_{\star}\Lambda_{c}|\rightarrow 0$, we demonstrated that the system can be described by atomic action given in Eq. 22. A small but non-zero $|\gamma|/|v_{\star}\Lambda_{c}|$ will generate coupling between moments formed by the $f$ electrons. In this section, we estimate the effective coupling in the projected limit. Since the flat bands are primarily composed of $f$ electrons with strong Hubbard interactions, we can apply a mean-field decoupling to the Hubbard repulsion term for the $f$ electrons.
$H_{U_{1}}\approx$ $-\frac{U_{1}}{2}\sum_{\mathbf{R}}\nu_{f}^{2}+U_{1}\nu_{f}\sum_{\mathbf{R},\alpha\eta s}:n_{\mathbf{R},\alpha\eta s}^{f}:$
$+\frac{U_{1}}{2}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R})O_{\alpha^{\prime}\eta^{\prime}s^{\prime},\alpha\eta s}(\mathbf{R})-U_{1}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R}):f_{\mathbf{R},\alpha^{\prime}\eta^{\prime}s^{\prime}}^{\dagger}f_{\mathbf{R},\alpha\eta s}:$ (49)
where
$O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R})=\langle:f_{\mathbf{R},\alpha\eta s}^{\dagger}f_{\mathbf{R},\alpha^{\prime}\eta^{\prime}s^{\prime}}:\rangle$
$\nu_{f}=\frac{1}{N}\sum_{\mathbf{R}}\text{Tr}[O(\mathbf{R})]$ (50)
and we have assumed the system does not develop CDW phase with a non-uniform charge distribution of $f$ electrons. We can observe that $O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R})$ describes the spin, valley, and orbital moments (flavor moments) formed by the $f$ electrons.
In terms of the path-integral, the corresponding action reads
$S=$ $S_{0}$
$+\int_{\tau}\bigg{[}\frac{U_{1}}{2}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R},\tau)O_{\alpha^{\prime}\eta^{\prime}s^{\prime},\alpha\eta s}(\mathbf{R},\tau)-U_{1}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R},\tau):f_{\mathbf{R},\alpha^{\prime}\eta^{\prime}s^{\prime}}^{\dagger}(\tau)f_{\mathbf{R},\alpha\eta s}(\tau):\bigg{]}$ (51)
where $S_{0}$ contains the single-particle term of the $f$ and $c$ electrons and can be written as
$S_{0}=$ $\frac{1}{\beta}\sum_{i\omega,\mathbf{k},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}\begin{bmatrix}f_{\mathbf{k},\alpha\eta s}^{\dagger}(i\omega)&c_{\mathbf{k},\alpha\eta s}^{\dagger}(i\omega)&c_{\mathbf{k},\alpha+3\eta s}^{\dagger}(i\omega)\end{bmatrix}$
$\begin{bmatrix}i\omega-\epsilon_{f}&\gamma&0\\ \gamma&i\omega-\epsilon_{c}&v_{\star}(\eta k_{x}+i(-1)^{\alpha+1}k_{y})\\ 0&v_{\star}(\eta k_{x}+i(-1)^{\alpha+1}k_{y})&i\omega-\epsilon_{c}\end{bmatrix}\begin{bmatrix}f_{\mathbf{k},\alpha\eta s}(i\omega)\\ c_{\mathbf{k},\alpha\eta s}(i\omega)\\ c_{\mathbf{k},\alpha+3\eta s}(i\omega)\end{bmatrix}-\frac{U_{1}}{2}\sum_{\mathbf{R}}\nu_{f}^{2}$ (52)
Here, $\epsilon_{f},\epsilon_{c}$ includes the Hartree contributions of $\hat{H}_{U_{1}},\hat{H}_{W},\hat{H}_{V}$ (see Ref. Hu et al. (2023b)).
We can then integrate out the $c$ and $f$ electrons which leads to the following action
$S_{eff}=$ $\int_{\tau}\bigg{[}\frac{U_{1}}{2}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R})O_{\alpha^{\prime}\eta^{\prime}s^{\prime},\alpha\eta s}(\mathbf{R})\bigg{]}-\text{Tr}\log\bigg{\{}[G_{0}^{-1}]-V[O]]\bigg{]}\bigg{\}}$ (53)
where $G_{0}$ is the Green’s function
$[G_{0}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=-\langle T_{\tau}f_{\mathbf{R},\alpha\eta s}(\tau)f^{\dagger}_{\mathbf{R}^{\prime},\alpha^{\prime}\eta^{\prime}s^{\prime}}(\tau^{\prime})\rangle$ (54)
calculated with respective to the action $S_{0}$, and the vertex that describes the coupling between $O(\mathbf{R})$ fields and electrons is defined as
$\bigg{[}V[O]\bigg{]}_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=\delta_{\mathbf{R}^{\prime},\mathbf{R}"}\delta_{\tau,\tau^{\prime}}O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R})U_{1}$ (55)
$S_{eff}$ can be understood as the effective action of the fields $O(\mathbf{R})$. From $S_{eff}$, we can derive the coupling between flavor moments ($O(\mathbf{R})$) of different sites.
We begin by discussing the properties of the Green’s function. Since we are focusing only on the contribution from the flat band, we could project the $f$ electron operator into the band basis ( Eq. 6)
$f^{\dagger}_{\mathbf{k},\alpha\eta s}(\tau)=\frac{1}{\sqrt{\mathcal{N}_{\mathbf{k}}}}d_{\mathbf{k},(-1)^{\alpha+1},\eta s}^{\dagger}+...$ (56)
where ... represents the contribution from the remote bands. To simplify the problem, we only include the contribution of the flat bands, since flat bands are near the Fermi energy. Then Green’s function could be written as
$[G_{0}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=$ $\frac{1}{N_{M}}\sum_{\mathbf{k}}\langle-T_{\tau}f_{\mathbf{k},\alpha\eta s}(\tau)f^{\dagger}_{\mathbf{k},\alpha^{\prime}\eta s^{\prime}}(\tau^{\prime})\rangle e^{i\mathbf{k}\cdot(\mathbf{R}-\mathbf{R}^{\prime})}$ (57)
Approximately, we find
$[G_{0}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}\approx$ $\frac{\delta_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}}{N_{M}}\sum_{\mathbf{k}}\frac{1}{{\mathcal{N}_{\mathbf{k}}}}\langle-T_{\tau}d_{\mathbf{k},(-1)^{\alpha+1},\eta s}(\tau)d^{\dagger}_{\mathbf{k},(-1)^{\alpha+1},\eta s}(\tau^{\prime})\rangle e^{i\mathbf{k}\cdot(\mathbf{R}-\mathbf{R}^{\prime})}$ (58)
where only the contributions from the flat band have been included.
To get a simple analytical expression, we assume the flat band is perfectly flat, meaning that $\langle d_{\mathbf{k},(-1)^{\alpha+1},\eta s}(\tau)d^{\dagger}_{\mathbf{k},(-1)^{\alpha+1},\eta s}(\tau^{\prime})\rangle$ is $\mathbf{k}$-independent. Under this assumption, we can introduce the following Green’s functions
[…]
For perfect flat bands at Fermi energy, the Green’s function takes the simple form of $g_{0}(i\omega)=1/i\omega$. Then we have
$[G_{0}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=$ $\frac{1}{\beta}\sum_{i\omega}\frac{1}{i\omega}e^{-i\omega(\tau-\tau^{\prime})}\frac{1}{N_{M}}\sum_{\mathbf{k}}\frac{1}{\mathcal{N}_{\mathbf{k}}}e^{i\mathbf{k}\cdot(\mathbf{R}-\mathbf{R}^{\prime})}\delta_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}$ (60)
where the $\mathbf{R}$-dependency of Green’s function arises from the wavefunction contribution, characterized by $1/\sqrt{\mathcal{N}_{\mathbf{k}}}$. For further convenience, we introduce
$I(\mathbf{R}-\mathbf{R}^{\prime})=\frac{1}{N_{M}}\sum_{\mathbf{k}}\frac{1}{\mathcal{N}_{\mathbf{k}}}e^{i\mathbf{k}\cdot(\mathbf{R}-\mathbf{R}^{\prime})}=\frac{1}{N_{M}}\sum_{\mathbf{k}}\frac{1}{1+\gamma^{2}/|v_{\star}\mathbf{k}|^{2}}e^{i\mathbf{k}\cdot(\mathbf{R}-\mathbf{R}^{\prime})}$ (61)
and then
$[G_{0}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=$ $\frac{1}{\beta}\sum_{i\omega}\frac{1}{i\omega}e^{-i\omega(\tau-\tau^{\prime})}I(\mathbf{R}-\mathbf{R}^{\prime})\delta_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}$ (62)
We are now in the position to evaluate the effective action defined in Eq. 53. We begin by separating the Green’s function into local and nonlocal contribution
$[G_{0}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=[G_{loc}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}+[\delta G]_{\tau\alpha\eta s,\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}$
$[G_{loc}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=\delta_{\mathbf{R},\mathbf{R}^{\prime}}[G_{0}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}$
$[\delta G]_{\tau\alpha\eta s,\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=(1-\delta_{\mathbf{R},\mathbf{R}^{\prime}})[G_{0}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}$ (63)
$G_{loc}$ and $\delta G$ correspond to the on-site (called here local) $(\mathbf{R}=\mathbf{R}^{\prime})$ and off-site (called here nonlocal) $(\mathbf{R}\neq\mathbf{R}^{\prime})$ contribution of Green’s function. The nonlocal Green’s function couples the flavor-moment fields $O(\mathbf{R})$ between different sites.
We then treat the nonlocal Green’s function $\delta G$ as a perturbation and derive the effective coupling of the flavor moments $O(\mathbf{R})$. With $G_{loc},\delta G$, we can rewrite $S_{eff}$ as
$S_{eff}=$ $\int_{\tau}\bigg{[}\frac{U_{1}}{2}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R},\tau)O_{\alpha^{\prime}\eta^{\prime}s^{\prime},\alpha\eta s}(\mathbf{R},\tau)\bigg{]}-\text{Tr}\log\bigg{\{}[G_{loc}+\delta G]^{-1}-V[O]\bigg{\}}$
$=$ $\int_{\tau}\bigg{[}\frac{U_{1}}{2}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R},\tau)O_{\alpha^{\prime}\eta^{\prime}s^{\prime},\alpha\eta s}(\mathbf{R},\tau)\bigg{]}$
$-\text{Tr}\log\bigg{\{}[G_{loc}+\delta G]^{-1}\bigg{\}}-\text{Tr}\log\bigg{\{}\mathbb{I}-[G_{loc}+\delta G]V[O]\bigg{\}}$ (64)
At zeroth order in $\delta G$, we have an atomic action
$S_{atom}\approx$ $\int_{\tau}\bigg{[}\frac{U_{1}}{2}\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}O_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(\mathbf{R},\tau)O_{\alpha^{\prime}\eta^{\prime}s^{\prime},\alpha\eta s}(\mathbf{R},\tau)\bigg{]}-\text{Tr}\log\bigg{\{}[G_{loc}]^{-1}-V[O]\bigg{\}}$ (65)
where each site is decoupled.
We can perform a Taylor expansion in powers of $\delta G$ which gives
$-\text{Tr}\log\bigg{\{}\mathbb{I}-[G_{loc}+\delta G]V[O]\bigg{\}}$
$=$ $\sum_{n=0}\frac{1}{n}\text{Tr}\bigg{\{}\bigg{[}(G_{loc}+\delta G)V[O]\bigg{]}^{n}\bigg{\}}$
$\approx$ $\sum_{n=0}\frac{1}{n}\text{Tr}\bigg{\{}\bigg{[}G_{loc}V[O]\bigg{]}^{n}\bigg{\}}+\sum_{n=2}^{\infty}\sum_{k=0}^{n-2}\text{Tr}\bigg{[}\delta GV[O]\bigg{(}G_{loc}V[O]\bigg{)}^{k}\delta GV[O]\bigg{(}G_{loc}V[O]\bigg{)}^{n-2-k}\bigg{]}\,.$ (66)
The leading-order coupling between flavor moments generated by the nonlocal Green’s function $\delta G$ is
$S^{\prime}=$ $\sum_{n=2}^{\infty}\sum_{k=0}^{n-2}\text{Tr}\bigg{[}\delta GV[O]\bigg{(}G_{loc}V[O]\bigg{)}^{k}\delta GV[O]\bigg{(}G_{loc}V[O]\bigg{)}^{n-2-k}\bigg{]}\ ,.$ (67)
We now evaluate $S^{\prime}$ explicitly. We first notice that the local Green’s function is
$[G_{loc}]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=\frac{1}{\beta}\sum_{i\omega}\delta_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}\delta_{\mathbf{R},\mathbf{R}^{\prime}}\frac{I(\bm{0})}{i\omega-\epsilon_{0}}e^{-i\omega(\tau-\tau^{\prime})}$ (68)
and the nonlocal Green’s function is
$[\delta G]_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=\frac{1}{\beta}\sum_{i\omega}\delta_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}(1-\delta_{\mathbf{R},\mathbf{R}^{\prime}})\frac{I(\mathbf{R}-\mathbf{R}^{\prime})}{i\omega-\epsilon_{0}}e^{-i\omega(\tau-\tau^{\prime})}$ (69)
As for the $V[O]$ term, we can introduce a $SU(8)$ rotation matrix $T(\mathbf{R})$ and a diagonal matrix $S$ such that
$O(\mathbf{R})=T(\mathbf{R})ST^{\dagger}(\mathbf{R})$ (70)
At integer filling, we approximately have $\nu_{f}\approx\nu\in\mathbb{Z}$. We can represent the diagonal matrix as
$S=\text{diag}\{\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2}\},\quad\nu=0$
$S=\text{diag}\{\frac{1}{2},\frac{1}{2},\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2}\},\quad\nu=-1$
$S=\text{diag}\{\frac{1}{2},\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2}\},\quad\nu=-2$
$S=\text{diag}\{\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2},-\frac{1}{2}\},\quad\nu=-3$ (71)
The diagonal components of $S$ with a value of $1/2$ correspond to the flavors that are filled, while those with a value of $-1/2$ correspond to the flavors that are empty. This gives $\text{Tr}[O(\mathbf{R})]=\text{Tr}[S]=\nu$. The decomposition in Eq. 70 gives
$\bigg{[}V[O]\bigg{]}_{\mathbf{R}\tau\alpha\eta s,\mathbf{R}^{\prime}\tau^{\prime}\alpha^{\prime}\eta^{\prime}s^{\prime}}=U_{1}\delta_{\mathbf{R},\mathbf{R}"}\delta_{\tau-\tau^{\prime}}\bigg{[}T(\mathbf{R})ST^{\dagger}(\mathbf{R})\bigg{]}_{\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}$ (72)
Combining Eqs. 67, 68, 69 and 72, we find
$S^{\prime}=\frac{1}{\beta}\sum_{i\omega,\mathbf{R},\mathbf{R}^{\prime}}\sum_{n=2}^{\infty}\sum_{k=0}^{n-2}\frac{I(\mathbf{R}-\mathbf{R}^{\prime})I(\mathbf{R}^{\prime}-\mathbf{R})I(\bm{0})^{n-2}}{(i\omega-\epsilon_{0})^{n}}U_{1}^{n}\text{Tr}\bigg{[}T_{\mathbf{R}^{\prime}}S^{1+k}T_{\mathbf{R}^{\prime}}^{\dagger}T_{\mathbf{R}}S^{n-1-k}T^{\dagger}_{\mathbf{R}}\bigg{]}$ (73)
Since $S$ is a diagonal matrix (Section IV.2), we find
$S^{n}=\frac{1}{2^{n-1}}S,\quad n\in 2\mathbb{Z}+1$
$S^{n}=\frac{1}{2^{n-1}}\mathbb{I},\quad n\in 2\mathbb{Z}$ (74)
This indicates that
$\text{Tr}\bigg{[}T_{\mathbf{R}^{\prime}}S^{n_{1}}T_{\mathbf{R}^{\prime}}^{\dagger}T_{\mathbf{R}}S^{n_{2}}T^{\dagger}_{\mathbf{R}}\bigg{]}=\frac{1}{2^{n_{1}+n_{2}-2}}8,\quad n_{1}\in 2\mathbb{Z},n_{2}\in 2\mathbb{Z}$
$\text{Tr}\bigg{[}T_{\mathbf{R}^{\prime}}S^{n_{1}}T_{\mathbf{R}^{\prime}}^{\dagger}T_{\mathbf{R}}S^{n_{2}}T^{\dagger}_{\mathbf{R}}\bigg{]}=\frac{1}{2^{n_{1}+n_{2}-1}}\text{Tr}[T_{\mathbf{R}^{\prime}}ST_{\mathbf{R}^{\prime}}^{\dagger}]=\frac{1}{2^{n_{1}+n_{2}-1}}\text{Tr}[O({\mathbf{R}^{\prime}})]=\frac{\nu_{f}}{2^{n_{1}+n_{2}-1}},\quad n_{1}\in 2\mathbb{Z}+1,n_{2}\in 2\mathbb{Z}$
$\text{Tr}\bigg{[}T_{\mathbf{R}^{\prime}}S^{n_{1}}T_{\mathbf{R}^{\prime}}^{\dagger}T_{\mathbf{R}}S^{n_{2}}T^{\dagger}_{\mathbf{R}}\bigg{]}=\frac{\nu_{f}}{2^{n_{1}+n_{2}-1}},\quad n_{1}\in 2\mathbb{Z},n_{2}\in 2\mathbb{Z}+1$
$\text{Tr}\bigg{[}T_{\mathbf{R}^{\prime}}S^{n_{1}}T_{\mathbf{R}^{\prime}}^{\dagger}T_{\mathbf{R}}S^{n_{2}}T^{\dagger}_{\mathbf{R}}\bigg{]}=\frac{1}{2^{n_{1}+n_{2}-2}}\text{Tr}[T_{\mathbf{R}^{\prime}}ST_{\mathbf{R}^{\prime}}^{\dagger}T_{\mathbf{R}^{\prime}}ST_{\mathbf{R}^{\prime}}^{\dagger}]=\frac{1}{2^{n_{1}+n_{2}-2}}\text{Tr}[O(\mathbf{R})O(\mathbf{R}^{\prime})],\quad n_{1}\in 2\mathbb{Z}+1,n_{2}\in 2\mathbb{Z}+1$ (75)
Only terms with $n_{1}\in 2\mathbb{Z}+1,n_{2}\in 2\mathbb{Z}+1$ generate a coupling between moments ($O(\mathbf{R})$) between different sites. Therefore, we focus on the terms with $11+k\in 2\mathbb{Z}+1,n-1-k\in 2\mathbb{Z}+$ 1 in  Eq. 73. This also implies that $n\in 2\mathbb{Z},k\in 2\mathbb{Z}$. We let $n=2m,k=2s$, and the corresponding contributions are
$S^{\prime}_{eff}=$ $\frac{1}{\beta}\sum_{i\omega,\mathbf{R},\mathbf{R}^{\prime}}\sum_{m=1}^{\infty}\sum_{s=0}^{m-1}\frac{I(\mathbf{R}-\mathbf{R}^{\prime})I(\mathbf{R}^{\prime}-\mathbf{R})I(\bm{0})^{2m-2}}{(i\omega-\epsilon_{0})^{n}}U_{1}^{2m}\text{Tr}\bigg{[}T_{\mathbf{R}^{\prime}}S^{1+2s}T_{\mathbf{R}^{\prime}}^{\dagger}T_{\mathbf{R}}S^{2m-1-2s}T^{\dagger}_{\mathbf{R}}\bigg{]}$
$=$ $\frac{1}{\beta}\sum_{i\omega,\mathbf{R},\mathbf{R}^{\prime}}\sum_{m=1}^{\infty}\sum_{s=0}^{m-1}\frac{I(\mathbf{R}-\mathbf{R}^{\prime})I(\mathbf{R}^{\prime}-\mathbf{R})I(\bm{0})^{2m-2}}{(i\omega-\epsilon_{0})^{2m}2^{2m-2}}U_{1}^{2m}\text{Tr}\bigg{[}O(\mathbf{R})O(\mathbf{R}^{\prime})\bigg{]}$
$=$ $\frac{1}{\beta}\sum_{i\omega,\mathbf{R},\mathbf{R}^{\prime}}I(\mathbf{R}-\mathbf{R}^{\prime})I(\mathbf{R}^{\prime}-\mathbf{R})\text{Tr}\bigg{[}O(\mathbf{R})O(\mathbf{R}^{\prime})\bigg{]}\frac{(i\omega)^{2}U^{2}}{\bigg{[}\omega^{2}+\bigg{(}\frac{I(\bm{0})U}{2}\bigg{)}^{2}\bigg{]}^{2}}$
$=$ $-\beta\sum_{\mathbf{R},\mathbf{R}^{\prime}}I(\mathbf{R}-\mathbf{R}^{\prime})I(\mathbf{R}^{\prime}-\mathbf{R})\frac{U}{4\pi I(\bm{0})}\text{Tr}[O(\mathbf{R})O(\mathbf{R}^{\prime})]$
$=$ $\beta\sum_{\mathbf{R},\mathbf{R}^{\prime}}J(\mathbf{R}-\mathbf{R}^{\prime})\text{Tr}[O(\mathbf{R})O(\mathbf{R}^{\prime})]$ (76)
In the last line, we have introduced the following coupling between the flavor moments
$J(\mathbf{R}-\mathbf{R}^{\prime})=-I(\mathbf{R}-\mathbf{R}^{\prime})I(\mathbf{R}^{\prime}-\mathbf{R})\frac{U}{4\pi I(\bm{0})}=-|I(\mathbf{R}-\mathbf{R}^{\prime})|^{2}\frac{U}{4\pi I(\bm{0})}$ (77)
Several comments are in order
  • $J(\mathbf{R}-\mathbf{R}^{\prime})$ is always ferromagnetic, which indicates the system always tends to develop ferromagnetic order in the limit we considered. We comment that a similar ferromagnetic coupling has also been derived in the unprojected limit of THF in Ref. Hu et al. (2023b).
  • We could refine our estimation of the effective spin-spin coupling by numerically evaluating Green’s function in Eq. 57, which will include the contributions from the dispersions of flat bands and the remote bands.
  • The coupling we derived is $U(8)$ symmetric. Additional symmetry-breaking terms in the THF model, which were ignored in the current calculation–such as the $v_{\star}^{\prime},M$–could generate effective couplings that break this $U(8)$ symmetry.
We can estimate the ordering temperatures of the ferromagnetic phase. The energy gained by developing a ferromagnetic order ($\mathbf{q}=0$ order) is proportional to
$J(\mathbf{q}=0)=\sum_{\mathbf{R}\neq\bm{0}}J(\mathbf{R})$ (78)
Here, we have subtracted $\mathbf{R}=0$ contribution, as the $\mathbf{R}=0$-term corresponds to a local coupling of the form $\text{[}O(\mathbf{R})O(\mathbf{R})]$ and does not favor any specific types of order. Written explicitly, we have
$J_{ferro}=J(\mathbf{q}=0)=$ $\sum_{\mathbf{R}}J(\mathbf{R})-J(\bm{0})=-\sum_{\mathbf{R}}|I(\mathbf{R})|^{2}\frac{U}{4\pi I(\bm{0})}+\frac{U}{4\pi}I(0)$
$=$ $-\frac{1}{N_{M}^{2}}\sum_{\mathbf{k}_{1},\mathbf{k}_{2}}\sum_{\mathbf{R}}\frac{U}{4\pi I(\bm{0})}\frac{1}{\mathcal{N}_{\mathbf{k}_{1}}\mathcal{N}_{\mathbf{k}_{2}}}e^{i(\mathbf{k}_{1}-\mathbf{k}_{2})\mathbf{R}}+\frac{U}{4\pi}I(0)$
$=$ $-\frac{1}{N_{M}}\sum_{\mathbf{k}_{1},\mathbf{k}_{2}}\frac{U}{4\pi}\frac{\delta_{\mathbf{k}_{1},\mathbf{k}_{2}}}{\mathcal{N}_{\mathbf{k}_{1}}\mathcal{N}_{\mathbf{k}_{2}}I(\bm{0})}+\frac{U}{4\pi}I(0)$ (79)
We then replace the $\mathbf{k}$ summation with a momentum integral
$\frac{1}{N_{M}}\sum_{\mathbf{k}}\approx\frac{1}{\Omega_{M}}\int_{|\mathbf{k}|<\Lambda_{c}}$ (80)
We can approximately take $\Omega_{M}\approx\pi\Lambda_{c}^{2}$ as we have discussed below Eq. 16. The momentum integral gives
$\frac{1}{N_{M}}\sum_{\mathbf{k}_{1}}\frac{1}{\mathcal{N}_{\mathbf{k}}^{2}}\approx$ $\frac{1}{\Omega_{M}}\int_{|\mathbf{k}|<\Lambda_{c}}\frac{1}{\bigg{[}1+\gamma^{2}/|v_{\star}\mathbf{k}|^{2}\bigg{]}^{2}}\approx 1+\frac{\gamma^{2}}{\Lambda_{c}^{2}|v_{\star}|^{2}}\bigg{(}1-\log\frac{\gamma^{4}}{v_{\star}^{4}\Lambda^{4}}\bigg{)}$ (81)
where we have also expanded the final results in powers of $\gamma^{2}/|\Lambda v_{\star}|^{2}$. In addition, we have
$I(\bm{0})=\frac{1}{N_{M}}\sum_{\mathbf{k}}\frac{1}{\mathcal{N}_{\mathbf{k}}}\approx\frac{1}{\Omega_{M}}\int_{|\mathbf{k}|<\Lambda_{c}}\frac{1}{1+\gamma^{2}/|v_{\star}\mathbf{k}|^{2}}=1-\frac{\gamma^{2}}{v_{\star}^{2}\Lambda_{c}^{2}}\log(1+\frac{v_{\star}^{2}\Lambda^{2}}{\gamma^{2}})$ (82)
Combining Sections IV.2, 81 and 82, we obtain
$J_{\rm ferro}\approx$ $-\frac{U}{\pi}\frac{\gamma^{2}}{v_{\star}^{2}\Lambda_{c}^{2}}\log\bigg{(}\frac{|\Lambda_{c}v_{\star}|}{|\gamma|}\bigg{)}$ (83)
We can then conclude the temperature scale for the development of ferromagnetic order is $T_{\rm ferro}\sim\frac{U}{\pi}\frac{\gamma^{2}}{v_{\star}^{2}\Lambda_{c}^{2}}\log\bigg{(}\frac{|\Lambda_{c}v_{\star}|}{|\gamma|}\bigg{)}$. We then expect the quasi-free flavor moments to appear within the temperature window $U_{1}\gtrsim T\gtrsim T_{\rm ferro}$. Finally, we comment that this ferromagnetic coupling persists in the regime where $U_{1}$ is larger than the gap between the remote band and flat band Hu et al. (2023b), allowing for continuity between the projected and the unprojected descriptions.
V Comparison between topological heavy-fermion model and non-local-moment model
In this section, we provide a detailed comparison between our results obtained from the THF model and those presented in Ref. Ledwith et al. (2024). As summarized in Table 1, the results in Ref. Ledwith et al. (2024) are consistent with our results. This consistency suggests that the THF model is applicable for studying both the strong coupling limit with $U_{1}>|\gamma|$ and the projected limit with $U_{1}<|\gamma|$. Furthermore, it suggests that the “non-local-moment behavior” discussed in Ref. Ledwith et al. (2024) could be understood via the Hubbard-I approximation of the THF model.
VI Quantitative aspects of the THF model
In this section, we discuss how the parameters of THF compare with experimental observations. In particular, $U_{1}$ is consistent with the cascade peaks in STM experiments as well as band gaps at fillings $\nu=\pm 4$. However, the agreement of other parameters with experiments remains less certain.
We adopt parameters at $\theta=1.05^{\circ}$ and $w_{0}/w_{1}=$0.8, 0.7. The corresponding parameters of the THF model, as presented in Ref. Song and Bernevig (2022) and Ref. Călugăru et al. (2023), are derived from the microscopic model, with a dielectric constant $\epsilon=6$ and a screening length $\xi=10$nm. Variations in the screening length and dielectric constant can alter these parameters. These parameters have also been used to study the local-moment behaviors in the DMFT calculations Hu et al. (2023b); Zhou et al. (2024); Rai et al. (2023). Here we plot the Hartree-Fock band structures (Fig. 2(a), (c)) at CNP with a Kramers inter-valley coherent order assumed. (As explained in Ref. Song and Bernevig (2022), the value of the gap does not depend on the order much.) The separation ${U}_{\rm sep}$ between two $f$-bands is found 50.6meV and 44.8meV for $w_{0}/w_{1}=$0.8 and 0.7, respectively. The DMFT ${U}_{\rm sep}$ for a symmetric state at CNP is about 40meV for $w_{0}/w_{1}=$0.8 Rai et al. (2023). ${U}_{\rm sep}$ is smaller than the bare $U_{1}$ parameter (58meV and 52meV) because of the $cf$ hybridization; and it is ${U}_{\rm sep}$ rather than $U_{1}$ should be compared to experiments. For comparison, we summarize ${U}_{\rm sep}$ seen in STM experiments of samples featuring quantum-dot-like cascades of transitions Wong et al. (2020b); Choi et al. (2021c); Oh et al. (2021b); Nuckolls et al. (2023b) that motivated the THF model. From Fig. 1d of Ref. Wong et al. (2020b), Fig. 4a of Ref. Choi et al. (2021c), Fig. 4g of Ref. Choi et al. (2021c), Fig. 1c of Ref. Oh et al. (2021b), Fig. 3a of Ref. Nuckolls et al. (2023b), one can extract $U_{\rm sep}\approx$45meV, 50meV, 49meV, 53meV, 40meV, respectively. As such, the theoretical $U_{\rm sep}$’s quantitatively match experimental values. Had it not matched the experiments, it would suggest an issue in the understanding of the parameters, since the values $U_{1}$ are not postulated but microscopically derived based on ab initio calculations and dielectric constants quoted in the experiment.
In addition, the compressibility obtained from the DMFT (Fig. 7a of Ref. Rai et al. (2023)) and Gutzwiller Hu et al. (2024) approaches both produce four cascades features from $\nu=0$ to $\nu=-4$, which is also consistent with experimental observation (Fig. 7a of Ref. Rai et al. (2023)). The changes of chemical potential from $\nu=0$ to $\nu=\pm 4$ obtained from DMFT calculations also match the experimental observations as shown in Fig. 7b of Ref. Rai et al. (2023).
We also calculate and plot the symmetric Hartree-Fock bands at $\nu=-4$ in Fig. 2(b), (d). The band gap is found 12.2meV and 25.2meV for $w_{0}/w_{1}=$0.8 and 0.7, respectively, consistent with the measured gap (15meV) by SET experiment Zondiner et al. (2020b).
Finally, we provide further discussions on the limit of
$\gamma/|v_{\star}\Lambda_{c}|\rightarrow 0\,.$ (84)
In this limit, the Berry curvature becomes highly concentrated and is effectively described by a $\delta$-function (Eq. 12). However, such a curvature can only be realized in a system with band touching. This implies that the gap between the remote bands and the dispersive bands vanishes, making it necessarily smaller than any finite $U_{1}$, which itself must be larger than the bandwidth of the narrow bands. This can be shown on general grounds (see Section VIII) and can also be observed within the THF model. If we fix the bandwidth of the $c$-electrons ($|v_{\star}\Lambda_{c}|$), then taking the limit $\gamma/|v_{\star}\Lambda_{c}|\to 0$ necessarily leads to $\gamma\to 0$, indicating that the gap between the remote and flat bands disappears in the heavy-fermion model. Thus, in this limit, due to band touching, one cannot simply focus on (or project onto) the flat bands while neglecting the remote bands. However, the heavy-fermion model, which explicitly incorporates both flat and remote bands, remains a valid framework for studying both regimes.
Figure 2: Hartree-Fock bands of the THF model. (a) and (c) are Hartree-Fock bands at CNP with parameters $w_{0}/w_{1}=0.8,0.7$, respectively. (b) and (d) are Hartree-Fock bands at $\nu=-4$ with parameters $w_{0}/w_{1}=0.8$, 0.7, respectively. The $U_{1}=$58.0, 51.7meV for $w_{0}/w_{1}=0.8$, 0.7, respectively.
VII Summary
In this work, we study the THF model of MATBG in the projected limit of
$|v_{\star}\Lambda_{c}|\gg|\gamma|\gg U_{1}$ (85)
We demonstrate that the THF model remains a faithful description of the system in the projected limit (Eq. 85), allowing for a straightforward analysis of the topology and correlations. In fact, the THF model was originally inspired by calculations in the momentum-space model within the projected limit Bernevig et al. (2021b); Lian et al. (2021); Bernevig et al. (2021c), where the single-particle excitation exhibits flat dispersion, indicating the presence of heavy $f$ electrons. Through the THF model, we have derived the flat-band wavefunction (also present in Ref. Song and Bernevig (2022)), the self-energy of the flat-bands, and the effective coupling between flavor moments.
Acknowledgements.
We thank D. Călugăru, T. Wehling, G. Sangiovanni, R. Valenti, P. J. Ledwith, and E. Khalaf for useful discussions. H. H. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101020833). H. H. was also supported by the Gordon and Betty Moore Foundation through Grant No.GBMF8685 towards the Princeton theory program, the Gordon and Betty Moore Foundation’s EPiQS Initiative (Grant No. GBMF11070), Office of Naval Research (ONR Grant No. N00014-20-1-2303), Global Collaborative Network Grant at Princeton University, BSF Israel US foundation No. 2018226, NSF-MERSEC (Grant No. MERSEC DMR 2011750), Simons Collaboration on New Frontiers in Superconductivity and the Schmidt Foundation at the Princeton University. Z.-D. S. was supported by National Natural Science Foundation of China (General Program No. 12274005), National Key Research and Development Program of China (No. 2021YFA1401900), and Innovation Program for Quantum Science and Technology (No. 2021ZD0302403). B. A. B was supported by DOE Grant No. DESC0016239.
VII.1 Appendix: Mapping an interacting Green’s function to a non-interacting Green’s function
In this appendix, we demonstrate that the Green’s function of an interacting system can be equivalently described by the Green’s function of non-interacting systems with additional auxiliary fermions. To show it, we start from a generic multi-orbital system with electron operators $\gamma_{i}$ where $i$ can denote momentum, orbital, sublattice indices. The system can be described by the following generic interacting Hamiltonian
$\hat{H}=\sum_{ij}t_{ij}\gamma_{i}^{\dagger}\gamma_{j}+\sum_{ijlm}U_{ijlm}\gamma_{i}^{\dagger}\gamma_{j}\gamma_{l}^{\dagger}\gamma_{m}$ (86)
where $t_{ij}$ is the hopping matrix, and $U_{ijlm}$ describes the interactions.
The Green’s function is then
$G_{ij}(\tau)=-\langle T_{\tau}\gamma_{i}(\tau)\gamma^{\dagger}_{j}(0)\rangle$
[…]
where $\Sigma_{ij}(i\omega)$ is the self-energy. We can utilize the spectral decomposition of the self-energy which reads
[…]
where $\rho^{\sigma}_{ij}(\epsilon)$ is the spectral functions of the self-energy functions and can be calculated from the real-frequency self-energy
$\rho_{ij}^{\Sigma}(\omega)=-\frac{1}{2\pi i}\left[\Sigma_{ij}(\omega+i0^{+})-\Sigma_{ji}^{*}(\omega+i0^{+})\right]$ (89)
We note that, from Eq. 89, the following properties of the self-energy holds
$\rho_{ij}^{\Sigma}(\omega)=\bigg{[}\rho_{ji}^{\Sigma}(\omega)\bigg{]}^{*}$ (90)
We then approximate the spectral functions $\rho_{ij}^{\Sigma}(\omega)$ via a series of $\delta(x)$ functions as
$\rho_{ij}^{\Sigma}(\omega)\approx\sum_{n}\bigg{(}v_{ij}^{n}\bigg{)}^{2}\delta(\omega-\epsilon_{ij}^{n})$ (91)
where $v_{ij}^{n},\epsilon_{n}$ are introduced to describe the behaviors of $\rho_{ij}^{\Sigma}$. From Eq. 89, we also note that
$v_{ij}^{n}=(v_{ji}^{n})^{*},\quad\quad\epsilon_{ij}^{n}=\epsilon_{ji}^{n}$ (92)
This also indicates that the self-energy can be written as
$\Sigma_{ij}(i\omega_{n})=\sum_{n}\frac{(v_{ij}^{n})^{2}}{i\omega_{n}-\epsilon_{ij}^{n}}$ (93)
We now prove that the interacting Green’s function (Eq. 87) is equivalent to the Green’s function of a non-interacting system.
We first introduce the following non-interacting systems
$\hat{H}^{aux}=\sum_{ij}t_{ij}\gamma_{i}^{\dagger}\gamma_{j}+\sum_{nij}\epsilon^{n}_{ij}a_{nij}^{\dagger}a_{nij}+\sum_{nij}v_{ij}^{n}\bigg{[}a_{nij}^{\dagger}\gamma_{i}+\gamma_{j}^{\dagger}a_{nij}\bigg{]}$ (94)
We can now calculate the corresponding Green’s functions with respect to the Hamiltonian $\hat{H}^{aux}$. The Green’s function (which is a matrix) is then
$G^{aux}(i\omega_{n})=\langle T_{\tau}\begin{bmatrix}\gamma(\tau)&a(\tau)\end{bmatrix}\cdot\begin{bmatrix}\gamma^{\dagger}(0)\\ a^{\dagger}(0)\end{bmatrix}\rangle$ (95)
where $\gamma(\tau)=[\gamma_{1}(\tau),...]$ is the vector formed by all the $\gamma$ electron operators, and $a(\tau)=[a_{111}^{\dagger}(\tau),...]$ is the vector formed by all the $a$ electron operators. The Green’s function then can be calculated as
$G^{aux}(i\omega_{n})=\rangle\begin{bmatrix}i\omega_{n}\mathbb{I}-t&-{V}\\ -V^{\dagger}&i\omega_{n}\mathbb{I}-\epsilon\end{bmatrix}^{-1}$ (96)
where the hybridization matrix is
$V_{i^{\prime},nji}=v_{ji}^{n}\delta_{i^{\prime},i},$ (97)
$t$ denotes the hopping matrix with element $t_{ij}$ for row $i$ and column $j$, and $\epsilon$ is a diagonal matrix with diagonal elements $\epsilon^{n}_{ij}$. We now aim to get the Green’s function of $\gamma$. We first note the following fact
$\begin{bmatrix}A&B\\ C&D\end{bmatrix}^{-1}=\begin{bmatrix}(A-BD^{-1}C)^{-1}&-(A-BD^{-1}C)^{-1}BD^{-1}\\ -(D-CA^{-1}B)^{-1}CA^{-1}&(D-CA^{-1}B)^{-1}\end{bmatrix}$ (98)
with $A,B,D,C$ matrices. Then we can calculate the Green’s function of $\gamma$ electrons by taking the upper left block of the $G^{aux}(i\omega_{n})$
$G^{aux,\gamma}_{ij}(i\omega_{n})=\langle T_{\tau}\gamma_{i}(\tau)\gamma_{j}(0)\rangle=\begin{bmatrix}i\omega_{n}\mathbb{I}-t-\tilde{\Sigma}(i\omega_{n})\end{bmatrix}^{-1}_{ij}$ (99)
where we have introduced
$\tilde{\Sigma}_{ij}(i\omega_{n})=\bigg{[}V\cdot[i\omega_{n}\mathbb{I}-\epsilon]^{-1}\cdot V\bigg{]}_{ij}=\sum_{n}\frac{v_{ij}^{n}v_{ij}^{n}}{i\omega_{n}-\epsilon_{ij}^{n}}$ (100)
By comparing Eq. 93 and Eq. 100, we can conclude that
$\tilde{\Sigma}_{ij}(i\omega_{n})=\Sigma_{ij}(i\omega_{n})$ (101)
Therefore, via Eq. 87 and Eq. 99, we conclude that
$G_{ij}^{aux,\gamma}(i\omega_{n})=G_{ij}(i\omega_{n})$ (102)
Then, we can conclude that, the interacting Green’s function can be equivalently described by the Green’s function of an effective non-interacting system. However, we note that, the mapping itself will not help us to calculate the interacting Green’s function, but allows us to more conveniently study the band topology of the interacting system via the effective non-interacting system.
VIII Appendix: Berry curvature and band gap in a two-orbital toy model
In this section, we analyze a simple two-orbital model to explore the relationship between the band gap and Berry curvature. We derive the Berry curvature at a general momentum $\mathbf{k}$ and show that a singular Berry curvature can only emerge when the band gap closes. Additionally, we examine the Berry curvature near the saddle point of the band and demonstrate that, as the gap closes, the Berry curvature can approach a $\delta$-function behavior near the saddle point.
We assume the hopping matrix of the two-orbital model (in two dimensions) takes the form of
$H_{\mathbf{k}}=\sum_{\mu=x,y,z}d_{\mu}(\mathbf{k})\sigma^{\mu}$ (103)
where $\sigma^{x,y,z}$ are Pauli matrices, and $d_{x,y,z}(\mathbf{k})$ are three real functions of $\mathbf{k}$. The dispersion reads
$E_{\mathbf{k},\pm}=\pm\sqrt{\sum_{\mu}[d_{\mu}(\mathbf{k})]^{2}}$ (104)
We assume the system is fully gapped with $\sqrt{\sum_{\mu}[d_{\mu}(\mathbf{k})]^{2}}\neq 0$ for all $\mathbf{k}$.
We now discuss the direct gap and the Berry curvature. For a generic $\mathbf{k}$ point $\mathbf{k}_{0}$, we could perform a small $k\cdot p$ expansion with
$H_{\mathbf{k}_{0}+(p_{x},p_{y})}=H_{\mathbf{k}_{0}}+p_{x}D_{x}[\mathbf{k}_{0}]+p_{y}D_{y}[\mathbf{k}_{0}]$ (105)
where $D_{x,y}[\mathbf{k}_{0}]$ are two $2\times 2$ Hermitian and tracless matrices (since $H_{\mathbf{k}}$ is traceless). We then perform a basis transformation denoted by $U(\mathbf{k}_{0})$ such that
$U(\mathbf{k}_{0})^{\dagger}H_{\mathbf{k}_{0}}U(\mathbf{k}_{0})=E_{\mathbf{k}_{0},+}\sigma_{z}$ (106)
We can further introduce
$h_{x}[\mathbf{k}_{0}]=U(\mathbf{k}_{0})^{\dagger}D_{x}[k_{0}]U(\mathbf{k}_{0}),\quad h_{y}[\mathbf{k}_{0}]=U(\mathbf{k}_{0})^{\dagger}D_{y}[k_{0}]U(\mathbf{k}_{0})$ (107)
Then the Hamiltonian near $\mathbf{k}_{0}$ can be written as
$U(\mathbf{k}_{0})^{\dagger}H_{\mathbf{k}_{0}+(p_{x},p_{y})}U(\mathbf{k}_{0})\approx E_{\mathbf{k}_{0}}\sigma_{z}+h_{x}[\mathbf{k}_{0}]p_{x}+h_{y}[\mathbf{k}_{0}]p_{y}$ (108)
In addition, we parametrize the $h_{x}[\mathbf{k}_{0}],h_{y}[\mathbf{k}_{0}]$ as
$h_{x}[\mathbf{k}_{0}]=v_{x}[\mathbf{k}_{0}][\mathbf{n_{x}}[\mathbf{k}_{0}]\cdot\bm{\sigma}],\quad h_{y}[\mathbf{k}_{0}]=v_{y}[\mathbf{k}_{0}][\mathbf{n_{y}}[\mathbf{k}_{0}]\cdot\bm{\sigma}]$ (109)
where $\mathbf{n_{x}}[\mathbf{k}_{0}],\mathbf{n_{y}}[\mathbf{k}_{0}]$ are two unit vectors, and
$v_{x/y}[\mathbf{k}_{0}]=\sqrt{\frac{1}{2}\text{Tr}[h_{x/y}[\mathbf{k}_{0}]\cdot h_{x/y}[\mathbf{k}_{0}]]}$ (110)
We now evaluate the Berry curvature. It is sufficient to consider the lowest band. The wavefunction can be written as
$u[\mathbf{k}_{0}+(p_{x},p_{y})]=\frac{1}{\sqrt{2\sqrt{X^{2}+Y^{2}+Z^{2}}(\sqrt{X^{2}+Y^{2}+Z^{2}}-Z})}\begin{bmatrix}Z-\sqrt{X^{2}+Y^{2}+Z^{2}}&X+iY\end{bmatrix},\quad$
$X=\sum_{\mu=x,y}p_{\mu}v_{\mu}[\mathbf{k}_{0}]n_{\mu}^{x},\quad X=\sum_{\mu=x,y}p_{\mu}v_{\mu}[\mathbf{k}_{0}]n_{\mu}^{y},\quad Z=\sum_{\mu=x,y}p_{\mu}v_{\mu}[\mathbf{k}_{0}]n_{\mu}^{z}+E_{\mathbf{k}_{0}}$ (111)
The Berry curvature can be calculated as
$\Omega(\mathbf{k}_{0}+(p_{x},p_{y}))=i\bigg{(}\partial_{x}u[\mathbf{k}_{0}+(p_{x},p_{y})]^{*}\partial_{y}u[\mathbf{k}_{0}+(p_{x},p_{y})]-\partial_{y}u[\mathbf{k}_{0}+(p_{x},p_{y})]^{*}\partial_{x}u[\mathbf{k}_{0}+(p_{x},p_{y})]\bigg{)}$ (112)
Exactly at $\mathbf{k}_{0}$, we observe the Berry curvature takes the form of
$\Omega(\mathbf{k}_{0})=\frac{1}{2}\frac{v_{x}[\mathbf{k}_{0}]v_{y}[\mathbf{k}_{0}]}{|E_{\mathbf{k}_{0},+}|^{2}}(\mathbf{n}_{x}[\mathbf{k}_{0}]\times\mathbf{n}_{y}[\mathbf{k}_{0}])\cdot\mathbf{e}_{z}$ (113)
Here, we could observe that $\Omega(\mathbf{k}_{0})$ is proportional to $1/|E_{\mathbf{k}_{0},+}|^{2}$. To obtain a diverged Berry curvature, we need to either let $|v_{x}[\mathbf{k}_{0}]v_{y}[\mathbf{k}_{0}|\rightarrow\infty$, or let $|E_{\mathbf{k}_{0},+}|^{2}\rightarrow 0$. However, we notice that (from Sections VIII, 110 and 105),
$v_{\mu}[\mathbf{k}_{0}]^{2}=\frac{1}{2}\text{Tr}\bigg{[}[\partial_{\mu}H_{\mathbf{k}_{0}}]^{2}\bigg{]}$ (114)
For a generic matrix with
$H_{\mathbf{k}}=\sum_{n,m,\mu}t_{n,m}^{\mu}\sigma^{\mu}e^{-i\mathbf{k}\cdot(n\mathbf{a}_{1}+m\mathbf{a}_{2})}$ (115)
where $t_{n,m}^{\mu}$ is the corresponding hopping strength, $\mathbf{a}_{1},\mathbf{a}_{2}$ are lattice vectors. We have
$[v_{\mu}[\mathbf{k}_{0}]]^{2}=\sum_{\nu}\bigg{|}\sum_{n,m}t_{n,m}^{\nu}(n\mathbf{a}_{1}+m\mathbf{a}_{2})_{\mu}\bigg{|}^{2}$ (116)
For an exponentially decayed hopping model, taking $|v_{\mu}[\mathbf{k}_{0}]|^{2}\rightarrow\infty$ is thus infeasible. Therefore, the only feasible way to realize a singular Berry curvature is to reduce the direct gap $|E_{\mathbf{k}_{0},+}|$ of the system at the corresponding momentum point. This then suggests a band-touching point.
Since the Berry curvature is proportional to $1/|E_{\mathbf{k}_{0},+}|^{2}$, it is also useful to investigate its behaviors near the saddle point, at which $1/|E_{\mathbf{k}_{0},+}|^{2}$ could reach its maximum/minimum. We consider a generic saddle point where
$\partial_{k^{\mu}}E_{\mathbf{k}_{0},\pm}=0$ (117)
Since
$E_{\mathbf{k}_{0}+(p_{x},p_{y}),\pm}$
$\approx$ ±$\pm\sqrt{(E_{\mathbf{k}_{0},+}+v_{x}[\mathbf{k}_{0}]{n}^{x}_{z}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{z}[\mathbf{k}_{0}]p_{y})^{2}+(v_{x}[\mathbf{k}_{0}]{n}^{x}_{x}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{x}[\mathbf{k}_{0}]p_{y})^{2}+(v_{x}[\mathbf{k}_{0}]{n}^{y}_{x}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{y}[\mathbf{k}_{0}]p_{y})^{2}}$
$\approx$ ±$\pm(E_{\mathbf{k}_{0},+}+v_{x}[\mathbf{k}_{0}]{n}^{x}_{z}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{z}[\mathbf{k}_{0}]p_{y})+O(p_{x}^{2},p_{y}^{2},p_{x}p_{y})$ (118)
Eq. 117 requires
$v_{x}[\mathbf{k}_{0}]\mathbf{n}_{x}[\mathbf{k}_{0}]\cdot\mathbf{e}_{z}=v_{y}[\mathbf{k}_{0}]\mathbf{n}_{y}[\mathbf{k}_{0}]\cdot\mathbf{e}_{z}=0$ (119)
This can be realized by taking $v_{x/y}[\mathbf{k}_{0}]=0$ or $\mathbf{n}_{x/y}[\mathbf{k}_{0}]\cdot\mathbf{e}_{z}=0$. If either $v_{x}[\mathbf{k}_{0}]=0$ or $v_{y}[\mathbf{k}_{0}]=0$, we have
$\Omega_{\mathbf{k}_{0}}=0$ (120)
which corresponds to a trivial saddle point. For the other situations with $\mathbf{n}_{x}[\mathbf{k}_{0}]\cdot\mathbf{e}_{z}=0$ and $\mathbf{n}_{x}[\mathbf{k}_{0}]\cdot\mathbf{e}_{z}=0$, we find
$\Omega(\mathbf{k}_{0}+(p_{x},p_{y}))=\frac{v_{x}[\mathbf{k}_{0}]v_{y}[\mathbf{k}_{0}]|E_{\mathbf{k}_{0},+}|(\mathbf{n}_{x}[\mathbf{k}_{0}]\times\mathbf{n}_{y}[\mathbf{k}_{0}])\cdot\mathbf{e}_{z}}{2\bigg{[}|E_{\mathbf{k}_{0},+}|^{2}+(v_{x}[\mathbf{k}_{0}]p_{x})^{2}+(v_{y}[\mathbf{k}_{0}]p_{y})^{2}+2v_{x}[\mathbf{k}_{0}]p_{x}v_{y}[\mathbf{k}_{0}]p_{y}(\mathbf{n}_{x}[\mathbf{k}_{0}]\cdot\mathbf{n}_{y}[\mathbf{k}_{0}])\bigg{]}^{3/2}}$ (121)
We could observe that
$\partial_{p_{\mu}}\Omega(\mathbf{k}_{0}+(p_{x},p_{y}))\bigg{|}_{p_{x}=p_{y}=0}=0$ (122)
which indicates that, the saddle point of the band dispersion also corresponds to the saddle point of the Berry curvature. The second-order derivative of the Berry curvature behaves as
$\partial_{p_{x}}^{2}\Omega(\mathbf{k}_{0}+(p_{x},p_{y}))\bigg{|}_{p_{x}=p_{y}=0}=-\frac{3(\mathbf{n}_{x}[\mathbf{k}_{0}]\times\mathbf{n}_{y}[\mathbf{k}_{0}])\cdot\mathbf{e}_{z}}{2|E_{\mathbf{k}_{0},+}|^{4}}(v_{x}[\mathbf{k}_{0}])^{3}v_{y}[\mathbf{k}_{0}]$
$\partial_{p_{y}}^{2}\Omega(\mathbf{k}_{0}+(p_{x},p_{y}))\bigg{|}_{p_{x}=p_{y}=0}=-\frac{3(\mathbf{n}_{x}[\mathbf{k}_{0}]\times\mathbf{n}_{y}[\mathbf{k}_{0}])\cdot\mathbf{e}_{z}}{2|E_{\mathbf{k}_{0},+}|^{4}}(v_{y}[\mathbf{k}_{0}])^{3}v_{x}[\mathbf{k}_{0}]$
$\partial_{p_{x}}\partial_{p_{y}}\Omega(\mathbf{k}_{0}+(p_{x},p_{y}))\bigg{|}_{p_{x}=p_{y}=0}=\frac{3(\mathbf{n}_{x}[\mathbf{k}_{0}]\cdot\mathbf{n}_{y}[\mathbf{k}_{0}])(\mathbf{n}_{x}[\mathbf{k}_{0}]\times\mathbf{n}_{y}[\mathbf{k}_{0}])\cdot\mathbf{e}_{z}}{2|E_{\mathbf{k}_{0},+}|^{4}}(v_{x}[\mathbf{k}_{0}])^{2}(v_{y}[\mathbf{k}_{0}])^{2}$ (123)
All the eigenvalues of matrix $M_{\mu\nu}=\partial_{p_{\mu}}\partial_{p_{\nu}}\Omega(\mathbf{k}_{0}+(p_{x},p_{y})))|_{p_{x}=p_{y}=0}$ have the same sign. This suggests that the Berry curvature reaches the local minimum/maximum correspondingly at the saddle point.
We next discuss the behaviors of the Berry curvature near $\mathbf{k}_{0}$. We could let
$p_{x}=p\cos(\theta),\quad p_{y}=p\sin(\theta)$
$v_{x}[\mathbf{k}_{0}]=V\cos(\theta_{v}),\quad v_{y}[\mathbf{k}_{0}]=V\sin(\theta_{v})$
$(\mathbf{n}_{x}[\mathbf{k}_{0}]\times\mathbf{n}_{y}[\mathbf{k}_{0}])\cdot\mathbf{e}_{z}=\sin(\phi),\quad\mathbf{n}_{x}[\mathbf{k}_{0}]\cdot\mathbf{n}_{y}[\mathbf{k}_{0}]=\cos(\phi)$ (124)
From Eq. 121, we find
$\Omega(\mathbf{k}_{0}+(p_{x},p_{y}))=\frac{V^{2}|E_{\mathbf{k}_{0},+}|\sin(\phi)\sin(2\theta_{v})}{4\bigg{[}|E_{\mathbf{k}_{0},+}|^{2}+V^{2}p^{2}f(\theta,\theta_{v},\phi)\bigg{]}^{3/2}}$
$f(\theta,\theta_{v},\phi)=\cos(\theta)^{2}\cos(\theta_{v})^{2}+\sin(\theta)^{2}\sin(\theta_{v})^{2}+\frac{1}{2}\sin(2\theta_{v})\sin(2\theta)\cos(\phi)$ (125)
We could observe that as we take $|E_{\mathbf{k}_{0}}|\rightarrow 0$ which indicates a band touching point, the Berry curvature distribution becomes a $\delta$ function.